### v1

# Leveraging Domain Knowledge for Efficient Reward Modelling in RLHF: A Case-Study in E-Commerce Opinion Summarization

Abstract
Reinforcement Learning from Human Feedback (RLHF) has become a dominating strategy in steering Language Models (LMs) towards human values/goals. The key to the strategy is employing a reward model ($\varphi$) which can reflect a latent reward model with humans. While this strategy has proven to be effective, the training methodology requires a lot of human preference annotation (usually of the order of tens of thousands) to train $\varphi$. Such large-scale preference annotations can be achievable if the reward model can be ubiquitously used. However, human values/goals are subjective and depend on the nature of the task. This poses a challenge in collecting diverse preferences for downstream applications. To address this, we propose a novel methodology to infuse domain knowledge into $\varphi$, which reduces the size of preference annotation required. We validate our approach in E-Commerce Opinion Summarization, with a significant reduction in dataset size (just $940$ samples) while advancing the state-of-the-art. Our contributions include a novel Reward Modelling technique, a new dataset (PromptOpinSumm) for Opinion Summarization, and a human preference dataset (OpinPref). The proposed methodology opens avenues for efficient RLHF, making it more adaptable to diverse applications with varying human values. We release the artifacts [^1] Code: github.com/efficient-rlhf. PromptOpinSumm: hf.co/prompt-opin-summ. OpinPref: hf.co/opin-pref for usage under MIT License.
Leveraging Domain Knowledge for Efficient Reward Modelling in RLHF: A Case-Study in E-Commerce Opinion Summarization
Swaroop ${}^{\dagger}$${}^{\clubsuit}$, Tejpalsingh ${}^{\dagger}$${}^{\clubsuit}$, Sri ${}^{\dagger}$${}^{\clubsuit}$, Rupasai ${}^{\dagger}$${}^{\clubsuit}$, Harshad ${}^{\diamondsuit}$, Pushpak ${}^{\clubsuit}$, Suman ${}^{\heartsuit}$, Amey ${}^{\heartsuit}$, Sudhanshu Shekhar ${}^{\heartsuit}$, Muthusamy ${}^{\heartsuit}$, Nikesh ${}^{\heartsuit}$$^{\clubsuit}$omputer Science and Engineering, IIT Bombay, India ${}^{\diamondsuit}$ Franklin Templeton ${}^{\heartsuit}$lipkart, India {swaroopnath, tejpalsingh, sriraghava, rupasai, pb}@cse.iitb.ac.in
${}^{\dagger}$ Equal contribution
1 Introduction
Reinforcement Learning from Human Feedback (RLHF) Ziegler et al. (2019); Ouyang et al. (2022) has emerged as a dominant paradigm in steering Language Models (LMs) towards human values. In the context of RLHF, human values are represented by a function ($\varphi$). For an output $Y$ ($=y_{1},y_{2},\cdots,y_{n}$) to some input $X$ ($=x_{1},x_{2},\cdots,x_{m}$), $\varphi$ performs the mapping $(X,Y)\rightarrow r$. The reward function $\varphi$ is latent to humans and manifests in human preferences. Preference Modelling techniques, such as Bradley-Terry model Bradley and Terry (1952), Plackett-Luce models Plackett (1975); Luce (2012) are used to learn $\varphi$ from preference data, of the form: $\mathcal{D}=\{(X,Y_{w},Y_{l})\>|\>Y_{w}\succ Y_{l}\}$.
Figure 1: Human Eval: Pairwise win-tie-loss percentage of Inductive-Bias model (our proposed model) vs. ground truth summary and summary from other models, for Amazon benchmark. We see that our proposed technique (domain knowledge infused reward model $\varphi$) helps Inductive-Bias model achieve summaries which are always preferred more than the competitors.
In contemporary works Ziegler et al. (2019); Bai et al. (2022a); Ouyang et al. (2022); Rafailov et al. (2023), the reward functions are parameterized as LMs themselves. Thus the raw text, ($X$, $Y_{w}$) and ($X$, $Y_{l}$) are directly fed to $\varphi$. Such a formulation necessitates large-scale human preference data, to train the large LM (typically millions/billions of parameters). Typically the size of $\mathcal{D}$ varies from $20K$ Ziegler et al. (2019); Nakano et al. (2021); Bai et al. (2022a) to $>200K$ Ethayarajh et al. (2022). Such a large-scale annotation might be achievable if the trained $\varphi$ can be used ubiquitously, irrespective of the nature of downstream application. However, human values are subjective in nature Jiang et al. (2022); Sorensen et al. (2023). For instance, hallucination would be desired in the task of Creative Writing, but undesired in Question-Answering. This means that depending on the context/downstream application, the reward function $\varphi$ can have varying characteristics. Collecting human preferences for all such downstream applications would be impractical.
Motivated to resolve this need, we propose a novel methodology for reward model training. We draw on the insight that $\varphi$ is dependent on the downstream application, and hence, can utilize its domain knowledge. Specifically, $\varphi$ lies in a low-dimensional manifold, whose dimensions can be deduced using domain knowledge. Such an inductive bias reduces the amount of samples [^2] An example: For a function, $f:(x_{1},x_{2},x_{3},\cdots,x_{m})\rightarrow y$, assuming that $f$ is a linear combination of $x_{i}$ (Linear Regression) reduces the training data requirement. Whereas, assuming no functional form (Feed-Forward Neural Network), would lead to more data requirement. needed to learn the parameters of $\varphi$, while increasing bias [^3] Bias of a Machine Learning model is the expected difference between its output and the true value. Say $f^{*}:X\rightarrow y$ is a function, and $\hat{f}$ is an estimate of $f^{*}$, from some training data, then bias is represented as: $\mathbb{E}[\hat{f}(X)-f^{*}(X)]$. of $\varphi$. Concretely our hypothesis is: An inductive bias infused $\varphi$ can help achieve alignment with human values for a task, with modest human preference annotations. We experimentally prove this in the domain of E-Commerce Opinion Summarization Bražinskas et al. (2020); Amplayo et al. (2021); Siledar et al. (2023b) – the task of summarizing user reviews for a product. First, we show that it helps advance the state-of-the-art in Opinion Summarization (Section 5). Second, we analyse how our approach helps the LM achieve alignment with human values for Opinion Summarization (Section 6).
Our contributions are:
  1. 1.
    A novel Reward Modelling technique for RLHF, which leverages Domain Knowledge, to achieve alignment with human values, while significantly reducing human preference annotation. In the domain of Opinion Summarization, we achieve alignment while reducing the dataset size by >21×>21$\times$> 21 × (as compared to the smallest publicly available preference data [^4] The smallest publicly available preference data is not in the domain of Opinion Summarization.). Our approach advances the state of the art: humans prefer our models’ outputs $>68\%$ over the SOTA.
  2. 2.
    A new dataset, called PromptOpinSumm, including reviews and summaries for $25763$ products ($229521$ summaries), to train and validate Opinion Summarizer models.
  3. 3.
    A new human preference dataset, called OpinPref, in the domain of Opinion Summarization, consisting of $940$ instances.
2 Related Works
Steering Language Models (LMs) towards human goals: Steering LMs towards human goals/values refers to the task of training LMs to generate text which is more aligned with human values, such as ‘text should not have harmful content’, ‘it should be polite’, etc. Such a task necessitates a human presence in the training loop of the these LMs. In recent times, Reinforcement Learning from Human Feedback (RLHF) Ziegler et al. (2019); Askell et al. (2021); Bai et al. (2022a); Ouyang et al. (2022); Liu et al. (2022) has emerged as a viable solution – by incorporating Reward Models, which reflect latent reward models wthin humans, into the training pipeline. These reward models are trained on human preference datasets Ziegler et al. (2019); Nakano et al. (2021); Ethayarajh et al. (2022), which are typically of the order of tens of thousands, in size. Dependence on high-quality large-sized preference data is an obstacle for RLHF.
Recently, Reinforcement Learning from AI Feedback (RLAIF) Bai et al. (2022b); Kim et al. (2023); Lee et al. (2023) has emerged as an alternative. It attempts to reduce the dependence on human preference datasets, by using Large LMs as preference data generators. While this is a scalable approach to steering LMs, there is no guarantee that the preference dataset generated by Large LMs actually reflects human goals – it is still a very open question. In our work, we propose a different solution – which promises to use human preference data, but provides a way to drastically reduce the required size. To the best of our knowledge, we are the first to attempt this.
Opinion Summarization: Opinion Summarization Hu and Liu (2004); Bražinskas et al. (2020); Amplayo et al. (2021); Siledar et al. (2023b) is the task of summarizing user reviews. Specifically, we look at E-Commerce Opinion Summarization, where user reviews are on products. These reviews contain aspects of the product, and users’ sentiments/opinions towards those aspects. Previous works Bražinskas et al. (2020); Siledar et al. (2023a) in E-Commerce Opinion Summarization have used Self-Supervised training methodology. In this context, self-supervision refers to picking one of the $N$ available reviews as a summary, commonly called pseudo-summary, and training the model on the remaining $N-1$ reviews to generate the pseudo-summary. The theme of solutions Chu and Liu (2018); Bražinskas et al. (2020); Siledar et al. (2023b, a) have mostly centered around Supervised Learning. The core problem has always been getting good synthetic datasets for training. More recently, Prompting Bhaskar et al. (2023) has been explored to solve the task. Bhaskar et al. (2023) move away from making a better synthetic dataset generation pipeline, and test Gpt-$3.5$ for Opinion Summarization.
We do not propose a new synthetic dataset generation methodology. Rather, we generate training data using an open-source Large LM (Mistral-$7$B), to test our hypothesis. To the best of our knowledge, we are the first to propose such a dataset for training Opinion Summarizers. Such an approach has been explored in earlier works Wang et al. (2023); Taori et al. (2023); Peng et al. (2023) for general domain text generation. Taori et al. (2023) fine-tune LLaMA-$7$B Touvron et al. (2023a) using Instruction-Tuning dataset generated using Gpt-$3$. Peng et al. (2023) fine-tune LLaMA-$7$B using a dataset generated by Gpt-$4$.
3 Dataset
Previous works Bražinskas et al. (2020); Siledar et al. (2023a) in Opinion Summarization have used Self-Supervised training methodology. In the context of Opinion Summarization, self-supervision refers to picking one of the $N$ available reviews as a summary, commonly called pseudo-summary, and training the model on the remaining $N-1$ reviews to generate the pseudo-summary. Among early works in this direction, Bražinskas et al. (2020) randomly selected a review as a pseudo-summary to construct a training sample. While, more recently, Siledar et al. (2023a) propose a sophisticated synthetic data creation pipeline for self-supervised opinion summarization.
Although these self-supervision datasets have helped further State-of-the-Art in Opinion Summarization, the trained models have several pitfalls. A few shortcomings are: (a) the generated opinion summaries are in first-person, (b) the review chosen as pseudo-summary might not cover all aspects and opinions, etc. To overcome these shortcomings, we move away from the self-supervised training regime in our work, and propose a new dataset. In the rest of this Section, we describe (a) PromptOpinSumm – a new dataset to train Opinion Summarization models, (b) the benchmarks we used to test our proposed technique, and (c) OpinPref – human preference annotated dataset for Opinion Summarization.
3.1 PromptOpinSumm Dataset
We prompt the Mistral-$7$B model Jiang et al. (2023) to generate opinion summary given reviews of a product. We also tried other open-source LLMs available at the time of the work, such as LLaMA2222-$7$B, LLaMA2222-$13$B Touvron et al. (2023b), Vicuna-$7$B, Vicuna-$13$B Chiang et al. (2023), Zephyr-$7$B Tunstall et al. (2023). However, we found that Mistral-$7$B leads to better summaries. We limit ourselves to open-source models mainly to limit the expenditure. Appendix E includes an examples along with qualitative analysis. We use the Amazon dataset He and McAuley (2016), which has reviews for $\sim 180k$ products. We sample reviews for $20763$ products for train set and $5000$ products for validation set to prompt the Mistral-$7$B model. Specifically, we prompt the model to generate opinion summaries of $3$ different qualities – Good (codenamed Good-Sum in the rest of the paper), Slightly Bad (codenamed SBad-Sum in the rest of the paper) and Very Bad (codenamed VBad-Sum in the rest of the paper). Again, per quality, we generate multiple opinion summaries ($3$ at most). We provide a reasoning of generating multiple summaries of different qualities in the discussion of our technique (Section 4.2). We generate a total of $184620$ summaries for train set and $44901$ summaries for validation set. Table 1 includes summary statistics of the generated dataset.
Split Characteristic $\mu$ $\sigma$
train # reviews $13.24$ $10.07$
# summaries $8.90$ $0.34$
# words in review $49.0$ $10.78$
# words in summary $78.28$ $34.45$
validation # reviews $10.53$ $6.80$
# summaries $8.98$ $0.16$
# words in review $48.65$ $10.63$
# words in summary $74.26$ $34.27$
Table 1: Statistics of PromptOpinSumm dataset.
3.2 Datasets for Evaluation
We test all our models on $3$ popular benchmarks: Amazon Bražinskas et al. (2020), Oposum+ Amplayo et al. (2021), and Flipkart Siledar et al. (2023b). Amazon dataset has reviews for $32$ products from $4$ domains: “electronics”, “home & kitchen”, “personal care”, and “clothing, shoes & jewellery”. Oposum+ dataset has reviews for $60$ products from $6$ domains: “laptop bags”, “bluetooth headsets”, “boots”, “keyboards”, “television”, and “vacuums”. Flipkart dataset has reviews for $147$ products from $3$ domains: “laptops”, “mobiles”, and “tablets”. Table 2 includes relevant statistics of these benchmarks.
Characteristic Oposum+ Amazon Flipkart
# domains $6$ $4$ $3$
# products $60$ $32$ 147
# reviews per product $10$ $8$ $10$
# summaries per product $3$ $3$ $1$
Table 2: Statistics of the benchmark datasets.
While we acknowledge the widespread use of these benchmarks, we notice several shortcomings. We highlight them in Appendix B. Due to these shortcomings, in our analysis, we do not rely much on overlap-based evaluations, such as Rouge. Rather we rely on human evaluations.
3.3 OpinPref Dataset
We create OpinPref by asking humans to rate opinion summaries for given reviews. We utilize domain experts (annotator details in Appendix G) to perform the annotation. We observe a Fleiss’ Kappa ($\kappa$) score of $62.67\%$ (substantial aggrement; aggrement is substantial when $60\%\leq\kappa<80\%$). We believe that aligning to the internal reward model of domain experts would lead to better opinion summaries. We provide the domain expert with reviews of a product and two opinion summaries (products are sampled from the PromptOpinSumm dataset). The domain expert notifies which of the two summaries they prefer. We use this to construct a dataset of the form: $\mathcal{D}_{h}=\{(R,s_{w},s_{l})\>|\>s_{w}\succ s_{l}\}$, where $R$ is the set of reviews and $s_{w}$ and $s_{l}$ are opinion summaries. We construct a dataset of $940$ samples. Appendix F includes statistics on the dataset (such as length of reviews and summaries).
Figure 2: Overview of our approach. Step-$1$: We generate a new dataset for training Opinion Summarizers: PromptOpinSumm, by prompting Mistral-$7$B  model. Again, we use Mistral-$7$B  to compute values for the $7$ features discussed in Section 4.1. Step-$2$: We ask humans (domain experts) to given their preference, given reviews and two opinion summaries (A, B). We use the preference data to train the reward model, $\varphi_{op}$. Step-$3$: We sample instances from PromptOpinSumm dataset; $\varphi_{op}$ assigns a score to the sampled summaries, the policy, $\pi_{\theta}$, assigns log probabilities to these summaries. Proximal Policy Optimization uses these to update $\pi_{\theta}$.
4 Technique
We test our hypothesis in the domain of Opinion Summarization in E-Commerce – the task of summarizing user reviews for a product. Typically, user reviews discuss several aspects of a product, opinions/sentiments towards these aspects. An opinion summary has to reflect all the aspects that the input reviews discuss, along with the opinions/sentiments expressed towards these aspects. The section is structured as follows: in Section 4.1, we discuss our reward modelling technique, human annotation, etc., and in Section 4.2, we discuss the RLHF training pipeline using the novel reward model.
4.1 Domain Knowledge Induced Reward Modelling
We leverage insights on desirable properties in an opinion summary from domain experts [^5] We consulted domain experts from an E-Commerce platform.. Based on these insights, we characterize our reward model, $\varphi_{op}$, as follows: $\varphi_{op}$$\>=f(v)$, where $v\in$aspect-coverage, opinion-faithfulness, opinion-coverage, conciseness, relevance, hallucination, language-correctness}. The features aspect-coverage, opinion-faithfulness  and opinion-coverage  are directly relevant to Opinion Summarization. They check if the generated opinion summary covers all mentioned aspects and opinions. The features conciseness, relevance  and hallucination  pertain to Generic Text Summarization Rush et al. (2015). They check if the generated summary is concise, relevant to the input reviews, and is free from hallucination. The features language-correctness  pertains to the grammar of the language, checking if the generated text follows the rules of the language. These features, together, characterize the goodness of an opinion summary. We use Mistral-$7$B (see Appendix A for details) to obtain values for these features for an opinion summary, given reviews. We denote this transformation using $\Phi$.
We train $\varphi_{op}$ using OpinPref, which is of the form: $\mathcal{D}_{h}=\{(R,s_{w},s_{l})\>|\>s_{w}\succ s_{l}\}$, where $R$ is the set of reviews and $s_{w}$ and $s_{l}$ are opinion summaries. We parameterize $\varphi_{op}$ using a Feed-Forward Neural Network, and train it using the Elo-loss Ouyang et al. (2022); Glaese et al. (2022) (Equation 4.1; $\Phi(R,s_{i})$ uses Mistral-$7$B to compute the $7$ features). Such a formulation for $\varphi_{op}$ brings interpretability – which features influence human preference the most, and is free from Alignment Tax Bai et al. (2022a) – degradation of LM on benchmarks after reward model training.
$\mathcal{L}_{pr}=-\mathbb{E}_{(R,s_{w},s_{l})\sim\mathcal{D}_{h}}\Big{[}$ $\log\sigma\big{(}\varphi_{op}(\Phi(R,s_{l}))$
$-\varphi_{op}(\Phi(R,s_{w}))\big{)}\Big{]}$ (1)
4.2 RLHF Training Pipeline
Using the trained reward model, we follow a similar training pipeline as Bai et al. (2022a); Ouyang et al. (2022), with a modification: Limited Trajectory Reinforcement Learning. Computing the transformation $\Phi$ for each generation online (during training) is expensive, especially with limited compute resources. To circumvent this, we limit the trajectories that are explored by our policy, $\pi_{\theta}$. Specifically, we limit it to the Good-Sum, SBad-Sum and VBad-Sum trajectories in the PromptOpinSumm dataset. Having varying levels of quality in PromptOpinSumm is of use here – it lets the model still explore trajectories of several quality. Thus, we have an offline experience buffer, with $\Phi$ precomputed, for $\pi_{\theta}$ learn from.
We use Proximal Policy Optimization (PPO) Schulman et al. (2017) to train our model (Equation 4.2). For each training step, we sample ($R,s,\Phi(R,s)$) tuples from PromptOpinSumm. We use the trained $\varphi_{op}$ to compute the reward for $s$ ($=\varphi_{op}\big{(}\Phi(R,s)\big{)}$). PPO uses this to update the log probability assigned by $\pi_{\theta}$. We parameterize $\pi_{\theta}$ using a Transformer model, which takes reviews as input, and generates an opinion summary.
$\mathcal{L}_{PPO}=-\mathbb{E}$ [${}_{(R,s,\Phi(s))}\bigg{[}\varphi_{op}(\Phi(R,s))$
$-\beta\log\Big{(}\dfrac{\pi^{RL}_{\theta}(s|R)}{\pi^{SFT}(s|R)}\Big{)}\bigg{]}$ (2)
Model-Code Amazon Flipkart Oposum+
R-$1\uparrow$ R-$2\>\uparrow$ R-L\>$\uparrow$ R-$1\>\uparrow$ R-$2\>\uparrow$ R-L\>$\uparrow$ R-$1\>\uparrow$ R-$2\>\uparrow$ R-L\>$\uparrow$
Prior Works MeanSum Chu and Liu (2018) $29.20$ $4.70$ $18.15$- -- -- $26.25$ $4.62$ $16.49$
CopyCat Bražinskas et al. (2020) $31.97$ $5.81$ $20.16$- -- -- $27.98$ $5.79$ $17.07$
PlanSum Amplayo and Lapata (2020) $32.87$ $6.12$ $19.05$- -- -- $30.26$ $5.29$ $17.48$
MultimodalSum Im et al. (2021) $34.19$ $7.05$ $20.81$- -- -- $33.08$ $7.46$ $19.75$
Op-Sum-GenSiledar et al. (2023a) $35.46$ $7.30$ $21.50$- -- -- $36.57$ $8.79$ $21.35$
Ours’ Supervised $28.22$ $4.91$ $16.79$ $28.10$ $4.18$ $15.37$ $30.15$ $7.26$ $16.74$
NaiveMean $28.09$ $4.89$ $16.75$ $26.42$ $4.14$ $15.0$ $30.62$ $7.62$ $17.11$
Synth-Feedback $24.82$ $4.30$ $15.48$ $27.40$ $3.64$ $15.25$ $29.10$ $6.19$ $16.20$
Inductive-Bias $29.16$ $4.74$ $17.07$ $27.14$ $4.0$ $15.0$ $30.75$ $7.45$ $16.82$
Table 3: Overlap-based Evaluation Results (R-$1$: Rouge-$1$, R-$2$: Rouge-$2$, R-L: Rouge-$L$). Our models are not near the state-of-the-art in these metrics. We expect such a trend, and we present the reasons in Section 3.2. We include these results for completeness. Note that for Flipkart benchmark, we do not have results from Op-Sum-Gen, as Siledar et al. (2023a) only provide aspect specific summarization model for the Flipkart benchmark.
5 Experiments and Results
Empirically, we test our technique exhaustively against the current State-of-the-Art model, and strong RL and RLHF baselines (our own design and from contemporary literature). We find that in overlap-based evaluation (Rouge-$1$, Rouge-$2$, Rouge-$L$), our proposed model (and other RL baselines) falls short of the SOTA. However, this is expected due to several shortcomings (Section 3.2) in the benchmark datasets. We conduct human and Gpt-$4$ evaluations to reliably verify efficacy of our technique. We find that our proposed technique excels significantly. In the rest of the section, we describe our models (Section 5.1) and evaluation results (Section 5.2).
5.1 Models
We train the following models on PromptOpinSumm:
Supervised: This is a supervised model trained using Maximum Likelihood Estimation.
NaiveMean: This is a Reinforcement Learning model, where the reward is computed by averaging the metric scores obtained using $\Phi$.
[…]
Inductive-Bias: This is a RLHF model, trained following our hypothesis. We train $\varphi_{op}$ using OpinPref dataset.
In addition to the models above, we also use the current SOTA Siledar et al. (2023a) for human evaluations. Siledar et al. (2023a) propose opinion summarizer models for general opinion summarization and aspect specific opinion summarization. As all of our models are for general opinion summarization, we use the same from Siledar et al. (2023a). We refer to this model by the name Op-Sum-Gen in the rest of the paper. Note that we do not use Direct Preference Optimization (DPO) Rafailov et al. (2023) and vanilla RLHF Ziegler et al. (2019); Bai et al. (2022a) as baselines. This is because both these techniques require huge human preference data. Additionally, the goal of the paper is not to propose a new RLHF technique, rather to propose a way to achieve alignment with domain-specific human goals with modest human preference annotation, using RLHF.
We use BART-Large Lewis et al. (2020) for all of our models. The choice of the model is governed by two factors: (a) It provides a similar environment (model size) for comparison with SOTA, (b) We find that LLMs (Mistral-$7$B, LLaMA2222-$7$B, Zephyr-$7$B, etc.) are already quite good at opinion summarization, thus any performance benefits (over SOTA) in those models cannot be reliably attributed to our novel reward modelling technique. We perform a hyperparameter sweep on batch size, learning rate, and learning rate warmup while training our models. We choose the model which performs best on the validation set of PromptOpinSumm to generate summaries for the benchmark datasets. We include more details on implementation in Appendix C.
5.2 Evaluation Results
On the models above, we perform overlap-based evaluations (Rouge). We report the scores in Table 3. In overlap-based evaluations, we find that our models do not perform better than existing models. We expect this behaviour (Section 3.2), which is why we do not draw conclusions from this evaluation. We report the metrics for completeness.
We also perform Gpt-$4$ and human evaluations. For human evaluations, we use domain experts (annotator details in Appendix G) for the evaluations; we observe a Fleiss’ Kappa ($\kappa$) score of $56.25\%$ (moderate agreement; agreement is moderate when $40\%\leq\kappa<60\%$). In both these settings, we ask the evaluator (Gpt-$4$ or domain expert) to rank the generated summaries given the reviews. We anonymize all the summaries and shuffle the orders for each evaluation instance. Given the rankings, we compute the fraction of pairwise wins, ties and losses among all the models. We compare summaries from Supervised, NaiveMean, Synth-Feedback, Inductive-Bias, Op-Sum-Gen models and ground truth summaries. We include ground truth summaries in the evaluation to verify our claims about the quality of the benchmarks. Figure 1 shows the performance of Inductive-Bias model over the others for human evaluations for the Amazon benchmark. We see that our proposed technique (domain knowledge based reward model $\varphi_{op}$) helps Inductive-Bias model achieve summaries which are always preferred more than the competitors. We can see a similar trend for Gpt-$4$ evaluation too, for all the benchmarks (Figures 3, 4 and 5). Both these evaluations empirically show that RLHF, using modest preference annotation, through domain knowledge infusion, can lead to performance gains.
Figure 3: Gpt-$4$ Eval: Pairwise win-tie-loss percentage of Inductive-Bias model vs. competitors, for Amazon benchmark.
Figure 4: Gpt-$4$ Eval: Pairwise win-tie-loss percentage of Inductive-Bias model vs. competitors, for the Flipkart benchmark. Note that for Flipkart benchmark, we do not have results from Op-Sum-Gen, as Siledar et al. (2023a) only provide aspect specific summarization model for Flipkart benchmark.
Figure 5: Gpt-$4$ Eval: Pairwise win-tie-loss percentage of Inductive-Bias model vs. competitors, for Oposum+ benchmark.
(a) aspect-coverage
(b) opinion-faithfulness
(c) relevance
(d) hallucination
Figure 6: Scores on domain knowledge based features on the Amazon benchmark for top-$3$ models (IB: Inductive-Bias, NM: NaiveMean, SF: Synth-Feedback). We see that Inductive-Bias model manages to stay ahead of the competitors, notably for hallucination. This increases our confidence on the Inductive-Bias model, as summaries generated by this model align more with the domain features.
6 Analysis
We perform a two-fold analysis: (a) First, we see the domain knowledge feature influences for $\varphi_{op}$, (b) Second, we see how the ground truth summary and summary from trained models fare on the domain knowledge features. This two-fold analysis helps us understand: (a) which features influence the latent reward model within humans [^6] Note that the trained $\varphi_{op}$ represents latent human reward model. the most, and (b) how the ground truth summary and summary from trained models fare on these influential features. Performing good on influential features would mean the summary aligns well with the latent reward model within humans.
6.1 Analysis of $\varphi_{op}$
$\varphi_{op}$ model has been trained on a set of features specified by domain experts. We analyze the relative influence of each feature on the final score assigned by $\varphi_{op}$. Doing this helps us understand an approximate importance [^7] We call this approximate importance as influence of a features on the output is not necessarily its importance. of each of these features. We do this by varying each feature by $\delta\>(=0.1)$ while keeping the other features constant, over multiple possible values of all features (Equation 6.1).
$\Delta_{i}=\dfrac{1}{2\delta}$ $\sum_{\mathbf{x}}\big{(}f(x_{1},\cdots,x_{i}+\delta,\cdots,x_{n})$
$-f(x_{1},\cdots,x_{i}-\delta,\cdots,x_{n})\big{)}$ (3)
Figure 7 highlights the relative influence of all the features. We see that hallucination is most influential. This aligns with what our human preference annotators report – hallucination in a summary is the primary cause of rejection. We see that, in majority cases, hallucinations are within the opinions in the summary. We see that this is also reflected in Figure 7opinion-faithfulness has significant influence. We also see that annotators prefer summaries which have more specifics about a product, i.e. they include more aspects – aspect-coverage has significant influence.
6.2 Analysis of Summaries
We analyze the top-$3$ performing models (in human and Gpt-$4$ evaluations) for the following features: opinion-coverage, opinion-faithfulness, hallucination and relevance. We show the analysis only for Amazon benchmark in the main body of the paper, we include rest in Appendix H. Figure 6 shows the performance on these features. We see that Inductive-Bias model fares much better than the competitors on hallucination (the most influential metric). For relevance, aspect-coverage and opinion-faithfulness, our model is fairly better than the other models.
Figure 7: Relative Influence of all features in $\varphi_{op}$. All the influences sum to $1$.
This shows that our technique helps Inductive-Bias model perform well on features that influence the latent reward model within humans for opinion summarization. This means that our technique helps Inductive-Bias model achieve a significant alignment with the latent reward model. This conclusion verifies our hypothesis (in the domain of opinion summarization) – An inductive bias infused reward model ($\varphi_{op}$) can help achieve alignment with latent reward model of humans for a task, with modest human preference annotations.
7 Summary, Conclusion and Future Work
In this work, we propose a novel Reward Modelling technique via Domain Knowledge Infusion. We verify our approach for E-Commerce Opinion Summarization, where we achieve state-of-the-art, while significantly reducing the amount of human preference annotations required (just $940$ samples). In addition to advancing SOTA and reducing preference annotations, our technique provides another two-fold benefits: (a) No Alignment Tax, and (b) Interpretability. Due to the interpretable nature, we find that our model does achieve alignment with human goals for Opinion Summarization through analysis. From the results and analysis, we conclude that Domain Knowledge Infusion into Reward Modelling is a viable solution to reduce human preference annotations for downstream tasks. In future, we would verify this for other domains.
8 Ethical Considerations
We contribute two datasets in our work: PromptOpinSumm, OpinPref. These datasets are generated using an open-source model Mistral-$7$B Jiang et al. (2023). We would release the datasets to further research in Opinion Summarization. For the OpinPref, we have seen that it does not contain any harmful content, such as social biases, stereotypes, etc. However, we have seen that it contains products of explicit nature (sexual products). For the PromptOpinSumm dataset, to the best of our knowledge, there is no presence of harmful content, such as social biases, stereotypes etc. We urge the research community to use the datasets with caution and check for potential harmfulness, based on their use-cases.
9 Limitations
A limitation of our work is we have tested our approach for one domain – Opinion Summarization. However, we do not believe that this weakens our argument, as we have exhaustively shown that our approach not only advances SOTA, but also interpretably achieves alignment with humans. Future work in other domains would help in verifying this claim for other domains. Another limitation is: we train our RLHF models using the limited trajectory trick. However, this limitation is imposed by the available compute resources. With larger compute resources, this study can be extended to extensive exploration during RLHF training.
Appendix A Features for Reward Modelling
We use $7$ domain specific features for the reward model $\varphi_{op}$. For each feature we prompt Mistral-$7$B to generate a score within 00 and $5$. For each feature, 00 means the model is doing bad on the feature, and $5$ means the model is doing good on the feature. We define all the features below:
aspect-coverage: This feature considers the aspect coverage within an opinion summary. The feature assumes a value $5$ if all the aspects of the product, mentioned in the reviews, are mentioned in the summary. If none of the aspects are picked, the feature assumes a value 00.
opinion-faithfulness: This feature considers whether the mentioned opinions/sentiments in the summary are correct, that is, they are picked correctly from the reviews. For example, if an user mentions that they are happy with the battery of a phone, and the summary mentions that users are unhappy with the battery, the summary will not be considered faithful to opinion in the review. The feature assumes a value $5$ if all the opinions are faithfully reflected. If no opinion is faithfully reflected, the value would be 00.
opinion-coverage: This feature considers whether all the opinions in the input reviews are picked by the opinion summary. The feature assumes a value $5$ if all the opinions are picked up. If none of the opinions are picked up, the feature assumes a value 00.
relevance: This feature checks if the summary is relevant to the input reviews (that is the product). The feature assumes a value $5$ if summary is completely relevant. If it is completely irrelevant, the feature assumes a value 00.
conciseness: This feature considers the conciseness and completeness of the opinion summary. The feature assumes a value $5$ if the summary is concise and complete – not one phrase/sentence can be dropped off. It assumes a value 00 if the summary is totally incomplete, or very verbose.
hallucination: This feature considers the factuality of the opinion summary. The feature assumes a value $5$ if the summary is totally factual, with respect to the input reviews. If there are a lot of hallucinations, the feature assumes a value 00.
language-correctness: This feature checks the correctness of language/text in the opinion summary. The feature assumes a value $5$ if the summary is grammatically fully correct. It assumes a value 00 if the summary is very poor linguistically.
For conciseness, we do not include the prompts in the paper, we would release them as separate artifacts, with the datasets, in the camera ready version.
Appendix B Example Summaries in Existing Benchmarks
We highlight the shortcomings of the benchmark datasets here.
Amazon: Bražinskas et al. (2020) designed the test-set in such a way that the summary has to read like a review, for instance, summary would contain ‘I think the quality has come down over the years.’, instead of ‘Users think that quality has come down over years’. Due to this writing style, the summaries read like reviews and are often in first person – high overlap would not necessarily mean a better summary, it would rather mean a better review.
Flipkart: Siledar et al. (2023b) generate this dataset by listing out the aspect-wise pros and cons presented within the reviews. We form an opinion summary by concatenating these pros and cons. Due to this, the summaries have frequent incoherent sentences.
Oposum+: Amplayo et al. (2021) create this benchmark by extracting sentences from the input reviews. Hence, this dataset has similar drawbacks as the Amazon benchmark.
Amazon
    • Nice boots but run a bit narrow. They look great but I think the quality has come down over the years. Still comfortable but I wish they broke in easier. I recommend these for any lady who is patient and looking for comfort.
Oposum+
    • great product for the cost . very easy to use and compatible with all of my phones ! it holds a charge great , is light enough and fits perfectly in my ear . the sound quality is great , the style is very cool and the unit feels top quality . it would drop and reconnect every 10 seconds nobody could hear me i could n’t get it to unpair from the phone , there ’s apparently no noise-cancellation in these . the battery life is … bizarre . cheap , plastic-y , and poor sound quality .
Flipkart
  • Summary
    • Pros
      • Design: The full-metal Infinix INBook X1 Core i3 has a top notch and premium design.
      • 35.56 cm (14 inch) 1920 x 1080 Pixel Full HD IPS Display: 100% sRGB with 300nits brightness ensures an excellent display.
      • Battery: Long-lasting battery. Gives around 8 hours of backup on normal usage.
      • Performance: The combination of Intel Core processor chip, high RAM size and sufficient storage capacity gives this laptop a high-speed performance experience.
      • Price: "Totally worth it in this price range.
    • Cons
      • Charging: Some current leakage during charging. Sometimes the laptop won’t charge.
      • Trackpad: Not upto the mark.
    • Verdict: This laptop comes with a i3 10th gen dual core processor which is suitable for normal tasks like web browsing, online classes and watching movies. Not recommended as a gaming laptop.
    • Additional Information: Can handle video editing and expandable SSD.
Appendix C Implementation Details
We use BART-Large Lewis et al. (2020) as our policy ($\pi_{\theta}$) in all of the models. We do this to have a fair comparison with the state-of-the-art in Opinion Summarization. We use AdamW Optimizer to train the models, with a weight decay of $0.05$. We use a cosine learning rate scheduler. We run a hyperparameter sweep on batch size, learning rate, and learning rate warmup. We include the possible values for the sweep in Table 4. We train all of our models using 2×2$\times$2 × A100100100100 GPUs ($80$GB)
Hyperparam Values
batch size [$64$, $128$, $256$]
learning rate $\sim\mathcal{U}(5e^{-6},5e^{-5})$
learning rate warmup $\sim\mathcal{U}(0.2,0.4)$
Table 4: Possible Values for Hyperparameters. For learning rate warmup, we sample the fraction of total steps the learning should be warmed up. For example, if the learning rate warmup is $0.2$, it means that the learning rate will have a linear warmup for $20\%$ of the total training steps.
For the reward model, $\varphi_{op}$, we use a Feed Forward Network for the Policy Model. We use AdamW Optimizer to train the models, with a weight decay of $0.05$. As before, we run a hyperparameter sweep on batch size, learning rate, and learning rate warmup. Table 5 includes details on the hyperparameters.
Hyperparam Values
batch size [$32$, $64$, $128$]
learning rate $\sim\mathcal{U}(5e^{-3},1e^{-1})$
Table 5: Possible Values for Hyperparameters for the Reward Model. For learning rate warmup, we sample the fraction of total steps the learning should be warmed up. For example, if the learning rate warmup is $0.2$, it means that the learning rate will have a linear warmup for $20\%$ of the total training steps.
Appendix D Generated Summary Lengths
We analyze the generation lengths of the models, and the ground truth summary. Table 6 lists the summary lengths.
Model Amazon Oposum Flipkart
Ground-Truth $60.65$ $85.86$ $129.91$
NaiveMean $79.93$ $99.76$ $71.59$
Synth-Feedback $87.59$ $115.00$ $84.18$
Op-Sum-Gen $55.84$ $62.93$
Inductive-Bias $73.53$ $68.63$ $68.63$
Supervised $81.84$ $126.00$ $77.80$
Table 6: Generation Length Statistics: number of words in summaries. We use NLTK to split into tokens.
Good: The users have shared their positive experiences with the protective covers for Macbooks. They appreciate the ease of application, vibrant colors, quick shipping, and great quality. Some have mentioned the covers fit their Macbooks perfectly, while others have shared their disappointment when they realized it did not fit their specific model. A few users have noted the covers add a personal touch and pizzaz to their devices. However, some users have reported issues with the covers not fitting their Macbooks or falling off, leaving them feeling disappointed and frustrated.
Slightly bad: Some users have reported positive experiences with the Macbook cover, praising its protective qualities, easy application, and vibrant colors. Others, however, have encountered issues with it not fitting properly on their devices or being returned due to size mismatches. Overall, the product has been described as cute, great quality, and worth the investment, though some buyers have experienced disappointment with its size compatibility and easy detachment.
Very Bad: This is a terrible product for people with old Macbooks, it doesn’t fit at all! And don’t even bother looking at the description carefully before buying.
Table 7: Example summaries of generated dataset
Appendix E Details of the Generated Dataset
We include an example from PromptOpinSumm dataset (Table 7). We show one sample from Good-Sum, SBad-Sum and VBad-Sum quality each. We do not include the reviews for conciseness. However, we incluce the salient aspects of the reviews. The reviews talk about the following things:
  1. 1.
    Great price, Nice looking / Good color, Good utility / Good protection, Quick shipping, Nice fitting, Good accessibility of the laptop while the cover is on, Good finish quality.
  2. 2.
    Not a good fit for older macbooks, Broken / Unusable for the original ($\sim 2009$ $2010$) white macbooks, cannot return return.
Good-Sum manages to discuss all of these things – it was able to detect that the cover does not fit specific models (highlighted in blue). It also detects that a few users like the “personal touch addition” factor of the cover.
VBad-Sum is also a good representation of a bad summary – it totally ignores the “positive” aspects of the product and presents the “negative” aspects only. It fails at Aspect Coverage, Opinion Faithfulness and Opinion Coverage.
SBad-Sum maintains almost a similar quality as the Good one. However, it fails to draw out certain aspects, such as “pizzaz”, “personal touch addition”, etc.
Appendix F Statistics of the OpinPref dataset
We look at the summary statistics for the OpinPref dataset. Table 8. We see that, interestingly, annotators prefer longer summaries – this is because these summaries contain more specifics and details from the reviews.
Characteristic Value
# words in reviews $641.21$
# reviews $13.08$
# Words in summaries $73.16$
# words in preferred summaries $85.41$
# words in unpreferred summaries $66.91$
Table 8: Statistics of the OpinPref dataset. We use NLTK to split into tokens.
Appendix G Annotator Details
We include two disjoint sets of annotators in our work – first for creation of OpinPref ($3$ annotators), second for human evaluation  ($3$ annotators). For both annotations, we use domain experts. The domain experts are NLP researchers (age group: $24-30$) who have worked in Opinion Summarization for a long time, with publication experience (in $*{}^{*}$ conferences). The domain experts for human evaluation also have a similar profile. The annotators have been paid generously, based on the standard annotation rates in the geographical location.
Appendix H All Evaluation Results
We include all evaluation results in this section. In Tables 9, 10, 11 and 11 we include pairwise comparison results, in a win/tie/loss format. We also include results on evaluation on the features in Tables 13, 14 and 15.
$\cdot$ Supervised NaiveMean Synth-Feedback Inductive-Bias Op-Sum-Gen
NaiveMean $0.50/0.06/0.38$
Synth-Feedback $0.44/0.12/0.44$ $0.40/0.09/0.5$
Inductive-Bias $\mathbf{0.56}/0.09/0.28$ $\mathbf{0.46}/0.18/0.31$ $0.56/0.12/0.28$
Op-Sum-Gen $0.31/0.28/0.38$ $0.25/0.12/0.56$ $0.25/0.21/0.5$ $0.25/0.06/\mathbf{0.68}$
Ground-Truth $0.46/0.06/0.48$ $0.31/0.18/0.44$ $0.40/0.15/0.40$ $0.28/0.09/\mathbf{0.59}$ $0.5/0.09/0.38$
Table 9: Pairwise Win/Tie/Loss Results for all models in Human Evaluation for Amazon Benchmark. We format the data as: win/tie/loss.
$\cdot$ Supervised NaiveMean Synth-Feedback Inductive-Bias Op-Sum-Gen
NaiveMean $0.63/0.12/0.25$
Synth-Feedback $0.59/0.12/0.28$ $0.5/0.06/0.44$
Inductive-Bias $\mathbf{0.62}/0.12/0.25$ $\mathbf{0.46}/0.09/0.44$ $\mathbf{0.5}/0.06/0.44$
Op-Sum-Gen $0.06/0.03/0.9$ 0.09/$0.0/0.9$ 0.0 / 0.90 $0.12/0.09/0.78$ $0.06/0.0/\mathbf{0.93}$
ground-truth $0.12/0.06/0.81$ $0.09/0.06/0.84$ $0.16/0.06/0.78$ $0.09/0.0/\mathbf{0.90}$ $0.68/0.09/0.22$
Table 10: Pairwise Win/Tie/Loss Results for all models in Gpt-$4$ Evaluation for Amazon Benchmark. We format the data as: win/tie/loss.
$\cdot$ Supervised NaiveMean Synth-Feedback Inductive-Bias
NaiveMean $0.57/0.12/0.30$
Synth-Feedback $0.57/0.06/0.36$ $0.52/0.12/0.36$
Inductive-Bias $\mathbf{0.63}/0.12/0.25$ $\mathbf{0.54}/0.16/0.30$ $\mathbf{0.57}/0.08/0.34$
Ground-Truth $0.10/0.06/0.84$ $0.06/0.01/0.92$ $0.07/0.01/0.91$ $0.06/0.02/\mathbf{0.91}$
Table 11: Pairwise Win/Tie/Loss Results for all models in Gpt-$4$ Evaluation for Flipkart Benchmark. We format the data as: win/tie/loss.
$\cdot$ Supervised NaiveMean Synth-Feedback Inductive-Bias Op-Sum-Gen
NaiveMean $0.56/0.03/0.4$
Synth-Feedback $0.5/0.16/0.34$ $0.46/0.1/0.44$
Inductive-Bias $\mathbf{0.66}/0.0/0.33$ $\mathbf{0.46}/0.1/0.44$ $\mathbf{0.56}/0.06/0.36$
Op-Sum-Gen $0.1/0.06/0.83$ $0.06/0.03/0.9$ $0.03/0.03/0.93$ $0.03/0.03/\mathbf{0.93}$
Ground-Truth $0.13/0.13/0.73$ $0.1/0.033/0.8666$ $0.06/0.06/0.86$ $0.06/0.06/\mathbf{0.86}$ $0.7/0.1/0.2$
Table 12: Pairwise Win/Tie/Loss Results for all models in Gpt-$4$ Evaluation for Oposum+ Benchmark. We format the data as: win/tie/loss.
$\cdot$ AC OPF OPC CC RL HL LC
Supervised $3.43\pm 0.20$ $3.71\pm 0.37$ $3.67\pm 0.26$ $3.79\pm 0.31$ $4.04\pm 0.37$ $3.89\pm 0.39$ $4.55\pm 0.35$
NaiveMean $3.56\pm 0.22$ $3.91\pm 0.50$ $3.76\pm 0.38$ $3.89\pm 0.36$ $4.04\pm 0.48$ $3.99\pm 0.48$ $4.60\pm 0.27$
Synth-Feedback $3.55\pm 0.40$ $3.87\pm 0.71$ $3.71\pm 0.43$ $3.94\pm 0.50$ $4.04\pm 0.61$ $3.94\pm 0.68$ $4.38\pm 0.92$
Inductive-Bias $\mathbf{3.60}\pm\mathbf{0.17}$ $\mathbf{3.95}\pm\mathbf{0.40}$ $\mathbf{3.85}\pm\mathbf{0.25}$ $3.99\pm 0.35$ $4.06\pm 0.34$ $\mathbf{4.07}\pm\mathbf{0.43}$ $\mathbf{4.65}\pm\mathbf{0.32}$
Op-Sum-Gen $3.34\pm 0.68$ $3.92\pm 0.79$ $3.70\pm 0.54$ $\mathbf{4.0}\pm\mathbf{0.50}$ $\mathbf{4.08}\pm\mathbf{0.72}$ $3.87\pm 1.08$ $4.05\pm 1.31$
Ground-Truth $3.55\pm 0.50$ $3.93\pm 0.46$ $3.56\pm 0.31$ $4.08\pm 0.32$ $4.04\pm 0.46$ $3.81\pm 0.86$ $4.40\pm 0.45$
Table 13: Intrisic Evaluation results on the Amazon Benchmark for all the models. Legend: AC: aspect-coverage, OPF: opinion-faithfulness, OPC: opinion-coverage, CC: conciseness, RL: relevance, HL: hallucination, LC: language-correctness.
$\cdot$ AC OPF OPC CC RL HL LC
Supervised $3.61\pm 0.22$ $4.10\pm 0.39$ $3.84\pm 0.33$ $4.04\pm 0.28$ $4.21\pm 0.31$ $4.19\pm 0.42$ $4.53\pm 0.27$
NaiveMean $3.56\pm 0.21$ $4.13\pm 0.41$ $3.84\pm 0.34$ $4.0\pm 0.32$ $\mathbf{4.31}\pm\mathbf{0.36}$ $4.26\pm 0.34$ $4.54\pm 0.39$
Synth-Feedback $3.56\pm 0.25$ $4.09\pm 0.40$ $3.79\pm 0.32$ $\mathbf{4.02}\pm\mathbf{0.30}$ $4.19\pm 0.34$ $4.19\pm 0.36$ $4.53\pm 0.29$
Inductive-Bias $\mathbf{3.63}\pm\mathbf{0.20}$ $\mathbf{4.22}\pm\mathbf{0.39}$ $\mathbf{3.85}\pm\mathbf{0.30}$ $4.01\pm 0.28$ $4.26\pm 0.29$ $\mathbf{4.33}\pm\mathbf{0.45}$ $\mathbf{4.61}\pm\mathbf{0.29}$
Op-Sum-Gen x x x x x x x
Ground-Truth $3.59\pm 0.15$ $3.88\pm 0.53$ $3.68\pm 0.27$ $4.02\pm 0.28$ $3.87\pm 0.59$ $3.67\pm 0.78$ $4.35\pm 0.44$
Table 14: Intrisic Evaluation results on the Flipkart Benchmark for all the models. Legend: AC: aspect-coverage, OPF: opinion-faithfulness, OPC: opinion-coverage, CC: conciseness, RL: relevance, HL: hallucination, LC: language-correctness.
$\cdot$ AC OPF OPC CC RL HL LC
Supervised $3.47\pm 0.14$ $3.38\pm 0.26$ $3.49\pm 0.06$ $\mathbf{3.64}\pm\mathbf{0.19}$ $\mathbf{3.81}\pm\mathbf{0.26}$ $3.22\pm 0.56$ $3.96\pm 0.32$
NaiveMean $3.49\pm 0.05$ $3.48\pm 0.06$ $3.5\pm 0.0$ $3.56\pm 0.13$ $3.66\pm 0.22$ $3.52\pm 0.33$ $4.1\pm 0.33$
Synth-Feedback $3.50\pm 0.03$ $3.41\pm 0.26$ $3.5\pm 0.0$ $3.63\pm 0.24$ $3.62\pm 0.20$ $3.32\pm 0.63$ $\mathbf{4.03}\pm\mathbf{0.38}$
Inductive-Bias $\mathbf{3.54}\pm\mathbf{0.22}$ $\mathbf{3.50}\pm\mathbf{0.06}$ $\mathbf{3.57}\pm\mathbf{0.06}$ $3.62\pm 0.19$ $3.65\pm 0.23$ $\mathbf{3.68}\pm\mathbf{0.36}$ $4.0\pm 0.29$
Op-Sum-Gen $3.39\pm 0.3$ $3.46\pm 0.45$ $3.49\pm 0.28$ $3.61\pm 0.40$ $3.58\pm 0.82$ $3.43\pm 0.92$ $3.79\pm 1.18$
Ground-Truth $3.42\pm 0.22$ $3.475\pm 0.28$ $3.5\pm 0.0$ $3.57\pm 0.16$ $3.49\pm 0.28$ $3.21\pm 0.48$ $3.56\pm 0.23$
Table 15: Intrisic Evaluation results on the Oposum+ Benchmark for all the models. Legend: AC: aspect-coverage, OPF: opinion-faithfulness, OPC: opinion-coverage, CC: conciseness, RL: relevance, HL: hallucination, LC: language-correctness.