### v2

# Leveraging Domain Knowledge for Efficient Reward Modeling in RLHF: A Case-Study in E-Commerce Opinion Summarization

Abstract
Reinforcement Learning from Human Feedback (RLHF) has become a dominating strategy in aligning Language Models (LMs) with human values/goals. The key to the strategy is learning a reward model ($\varphi$), which can reflect the latent reward model of humans. While this strategy has proven effective, the training methodology requires a lot of human preference annotation (usually in the order of tens of thousands) to train $\varphi$. Such a large-scale annotation is justifiable when it’s a one-time effort, and the reward model is universally applicable. However, human goals are subjective and depend on the task, requiring task-specific preference annotations, which can be impractical to fulfill. To address this challenge, we propose a novel approach to infuse domain knowledge into $\varphi$, which reduces the amount of preference annotation required (21×21$\times$21 ×), omits Alignment Tax, and provides some interpretability. We validate our approach in E-Commerce Opinion Summarization, with a significant reduction in dataset size (to just $940$ samples) while advancing the SOTA ($\sim 4$ point Rouge-$L$ improvement, $68\%$ of times preferred by humans over SOTA). Our contributions include a novel Reward Modeling technique and two new datasets: PromptOpinSumm (supervised data for Opinion Summarization) and OpinPref (a gold-standard human preference dataset). The proposed methodology opens up avenues for efficient RLHF, making it more adaptable to applications with varying human values. We release the artifacts [^1] Code: github.com/efficient-rlhf. PromptOpinSumm: hf.co/prompt-opin-summ. OpinPref: hf.co/opin-pref for usage under MIT License.
Leveraging Domain Knowledge for Efficient Reward Modeling in RLHF: A Case-Study in E-Commerce Opinion Summarization
Swaroop Nath, Tejpalsingh Siledar, Sri Raghava, Rupasai Rangaraju, Harshad Khadilkar, Pushpak Bhattacharyya, Suman Banerjee, Amey Patil, Sudhanshu Shekhar Singh, Muthusamy Chelliah, Nikesh Garera Computer Science and Engineering, IIT Bombay, India Flipkart, India swaroopnath@cse.iitb.ac.in
Equal contribution
1 Introduction
[…]
Figure 1: Human Eval: Pairwise win-tie-loss percentage of Inductive-Bias model (our proposed model) vs. ground truth summary and summary from other models, for Amazon benchmark. We see that our proposed approach (infusing domain knowledge into $\varphi$ to reap benefits of RLHF with modest human preference data) helps Inductive-Bias model achieve summaries which are always preferred (Section 5.2).
In contemporary works Ziegler et al. (2019); Bai et al. (2022a); Ouyang et al. (2022); Rafailov et al. (2023), the reward functions are Large LMs (LLMs; pretrained Transformers) themselves. The text data, ($X$, $Y_{w}$) and ($X$, $Y_{l}$) are directly fed to $\varphi$, for training. Such a formulation necessitates large-scale human preference data to train the LLM (millions/billions of parameters). Typically the size of $\mathcal{D}$ varies from $20K$ Nakano et al. (2021); Bai et al. (2022a) to $>200K$ Ethayarajh et al. (2022). Such a large-scale annotation is justifiable when it’s a one-time effort, and the trained $\varphi$ is universally applicable, irrespective of the nature of the downstream task. However, human values are subjective Jiang et al. (2022); Sorensen et al. (2023). For instance, hallucination would be desired in Creative Writing, but not in Question-Answering. This means that depending on the downstream task, the reward function $\mathbf{\varphi}$ must have varying characteristics. Collecting human preferences for all such tasks is impractical.
Motivated to resolve this need, we propose a novel reward modeling methodology, significantly reducing preference data requirements. We draw on the insight that $\varphi$ is dependent on the downstream task and, hence, can utilize its task/domain [^3] We use task and domain interchangeably in the paper. knowledge. Specifically, $\varphi$ lies in a low-dimensional manifold, whose dimensions can be deduced using domain knowledge. Such an inductive bias reduces the number of samples [^4] An example: For a function, $f:(x_{1},x_{2},x_{3},\cdots,x_{m})\rightarrow y$, assuming that $f$ is a linear combination of $x_{i}$ (Linear Regression) reduces the training data requirement. Assuming no functional form (Feed-Forward Neural Network) would require more data. needed to train $\varphi$. Concretely, our hypothesis is: An inductive bias infused $\varphi$ can help achieve alignment with human values for a task, with modest human preference annotations. Specifically, we say that $\varphi_{\tau}$ (reward model for a domain $\tau$) can be modelled by some numeric features $v_{1},v_{2},\cdots,v_{n}$. These $n$ features fully characterize [^5] Example of such characterization: Features like fluency, coherence, etc. can characterize text generated by an LLM. the outputs from the LLM on some input. Thus, instead of training $\varphi_{\tau}$ on the text data ($\{(X,Y_{w},Y_{l})\>|\>Y_{w}\succ Y_{l}\}$), we use the $n$ features. Such a formulation for $\varphi$ brings interpretability—which features influence human preference the most (Section 6), and is free from Alignment Tax (degradation of language capabilities of an LLM post reward modeling; Bai et al. (2022a)) as we do not use an LLM to model $\varphi$.
We experimentally prove our hypothesis in the domain of E-Commerce Opinion Summarization Bražinskas et al. (2020); Amplayo et al. (2021); Siledar et al. (2023b)—the task of summarizing user reviews for a product. In addition to advancing SOTA, we also analyze how our approach helps the model achieve alignment with human values for Opinion Summarization (Section 6).
Our contributions are:
  1. 1.
    A novel Reward Modeling technique for RLHF, which leverages Domain Knowledge to achieve alignment with human values while significantly reducing human preference annotation. In the domain of Opinion Summarization, we achieve alignment while reducing [^6] As compared to the smallest publicly available preference data. The smallest publicly available preference data is not in the domain of Opinion Summarization. the dataset size by $>\mathbf{21\times}$. Our approach advances SOTA: at least $\sim\mathbf{4}$-point Rouge-$L$ improvement (Tables 1, 4 and 5; Section 5.2), and humans prefer our models’ outputs $\mathbf{68}\%$ over SOTA (Figure 1; Section 5.2).
  2. 2.
    Two new datasets: PromptOpinSumm and OpinPref. PromptOpinSumm includes reviews and summaries for $25763$ products ($229521$ summaries), for training and validation. OpinPref is a gold-standard human preference dataset (with $940$ instances) in the domain of Opinion Summarization.
2 Related Works
Steering Language Models (LMs) towards human goals: Steering LMs towards human goals/values refers to the task of training LMs to generate text which is more aligned with human values, such as ‘text should not have harmful content’, ‘it should be polite’, etc. Such a task necessitates a human presence in the training loop of these LMs. In recent times, Reinforcement Learning from Human Feedback (RLHF) Ziegler et al. (2019); Askell et al. (2021); Bai et al. (2022a); Ouyang et al. (2022); Liu et al. (2022) has emerged as an effective solution—by incorporating Reward Models, which reflect latent reward models within humans, into the training pipeline. These reward models are trained on human preference datasets Ziegler et al. (2019); Nakano et al. (2021); Ethayarajh et al. (2022), which are typically of the order of tens of thousands, in size. Dependence on high-quality, large-sized preference data is an obstacle for RLHF.
Recently, Reinforcement Learning from AI Feedback (RLAIF) Bai et al. (2022b); Kim et al. (2023); Lee et al. (2023) has emerged as an alternative. It attempts to reduce the dependence on human preference datasets by using Large LMs (LLMs) as preference data generators. While this is a scalable approach to steering LMs, there is no guarantee that the preference dataset generated by LLMs reflects human goals. In our work, we propose a different solution, which promises to use human preference data but provides a way to reduce the required size drastically. To the best of our knowledge, we are the first to attempt this.
Opinion Summarization: Opinion Summarization Hu and Liu (2004); Bražinskas et al. (2020); Amplayo et al. (2021); Siledar et al. (2023b) is the task of summarizing user reviews. Specifically, we look at E-Commerce Opinion Summarization, where user reviews are on products. These reviews contain aspects of the product and users’ sentiments/opinions towards those aspects. Previous works Bražinskas et al. (2020); Siledar et al. (2023a) in E-Commerce Opinion Summarization have used Self-Supervised training methodology. In this context, self-supervision refers to picking one of the $N$ available reviews as a summary, commonly called pseudo-summary, and training the model on the remaining $N-1$ reviews to generate the pseudo-summary. The theme of solutions Chu and Liu (2018); Bražinskas et al. (2020); Siledar et al. (2023b, a) have mostly centered around Supervised Learning. The core problem has always been getting good synthetic datasets for training. More recently, Prompting Bhaskar et al. (2023) has been explored to solve the task. Bhaskar et al. (2023) move away from making a better synthetic dataset generation pipeline and test Gpt-$3.5$ for Opinion Summarization.
We do not propose a new synthetic dataset generation methodology. Rather, we generate training data using an open-source LLM (Mistral-$7$B), to test our hypothesis. To the best of our knowledge, we are the first to propose such a dataset for training Opinion Summarizers. Such an approach has been explored for Generic Text Summarization Wang et al. (2023); Taori et al. (2023); Peng et al. (2023). Taori et al. (2023) fine-tune LLaMA-$7$B Touvron et al. (2023a) using Instruction-Tuning dataset generated using Gpt-$3$. Peng et al. (2023) fine-tune LLaMA-$7$B using a dataset generated by Gpt-$4$.
3 Dataset
Previous works Bražinskas et al. (2020); Siledar et al. (2023a) in Opinion Summarization have used Self-Supervised training methodology, where $N-1$ reviews are used as input, and the left out review is used as a pseudo-summary (Section 2). Although these self-supervision datasets have helped further Opinion Summarization research, the approach has several shortcomings: the summaries always present a one-person rather than the consensus view, the summaries are reviews and might not have good coverage of aspects and opinions, etc. We move away from self-supervision to overcome these shortcomings and propose a new dataset. In the rest of this Section, we describe (a) PromptOpinSumm: a new dataset to train Opinion Summarizers, (b) the benchmarks we used for evaluation, and (c) OpinPref: gold-standard preference dataset for Opinion Summarization.
3.1 PromptOpinSumm Dataset
We prompt the instruction-tuned Mistral-$7$B model Jiang et al. (2023) to generate an opinion summary given product reviews. We also tried other open-source LLMs available at the time of the work, such as LLaMA2222-$7$B, LLaMA2222-$13$B Touvron et al. (2023b), Vicuna-$7$B, Vicuna-$13$B Chiang et al. (2023), Zephyr-$7$B Tunstall et al. (2023). However, we found that Mistral-$7$B leads to better summaries. We limit ourselves to open-source models due to cost. Appendix G includes examples and qualitative analysis. We use the Amazon dataset He and McAuley (2016), which has reviews for $\sim 180k$ products. We randomly sample reviews for $20763$ products for train set and $5000$ products for validation set. Specifically, we prompt the model to generate opinion summaries of $3$ different qualities: Good (codenamed Good-Sum), Slightly Bad (codenamed SBad-Sum), and Very Bad (codenamed VBad-Sum). We generate multiple opinion summaries ($3$ at most) per quality. We provide reasoning for generating multiple summaries of different qualities in the extended discussion of our approach (Appendix B). We generate $184620$ summaries for train set and $44901$ summaries for validation set (see Appendix G).
3.2 Benchmarks for Evaluation
We use $9$ Opinion Summarization benchmarks for evaluation. $3$ of these benchmarks are the Amazon test set (Bražinskas et al. (2020), codenamed Amazon), the Oposum+ test set (Amplayo et al. (2021), codenamed Oposum+) and the Flipkart test set (Siledar et al. (2023b), codenamed Flipkart). Amazon has reviews for $32$ products from $4$ domains, Oposum+ has reviews for $60$ products from $6$ domains and Flipkart has reviews for $147$ products from $3$ domains.
Although these $3$ benchmarks have been used widely, they have several shortcomings. For instance, Amazon was developed by asking annotators to write a summary in first-person point-of-view. This causes problems such as summaries seeming personal rather than consensus opinions (which can include mixed sentiment), incomplete coverage of aspects and opinions, etc. Thus, using such pseudo-summaries for reference-based evaluations (Rouge, BertScore) on such a benchmark is not a correct portrayal of the models’ performances. We highlight the shortcomings in detail in Appendix D. Siledar et al. (2024) recently provided $6$ new benchmarks (Amazon-R, Amazon-Rdq, Oposum-R, Oposum-Rdq, Flipkart-R, Flipkart-Rdq) which are revamped versions (by getting rid of the shortcomings) of the aforementioned $3$ benchmarks. We primarily rely on these $6$ for our conclusions. Appendix D includes more details on domains and summary statistics.
3.3 OpinPref Dataset
We create OpinPref by asking humans to rank opinion summaries for given reviews. We utilize domain experts (annotator details in Appendix I) to perform the annotation. We believe that aligning with the internal reward model of domain experts would lead to better opinion summaries. We provide the domain expert with product reviews and two opinion summaries (products are sampled from the PromptOpinSumm dataset). The domain expert notifies which of the two summaries they prefer. We use this to construct a dataset of the form: $\mathcal{D}_{h}=\{(R,s_{w},s_{l})\>|\>s_{w}\succ s_{l}\}$, where $R$ is the set of reviews and $s_{w}$ and $s_{l}$ are opinion summaries. We construct a dataset of $940$ samples. We observe a Fleiss’ Kappa ($\kappa$) score of $62.67\%$ (substantial aggrement; aggrement is substantial when $60\%\leq\kappa<80\%$). Appendix H includes statistics on the dataset.
Figure 2: Overview of our approach. Step-$1$: We generate a new dataset for training Opinion Summarizers: PromptOpinSumm, by prompting Mistral-$7$B  model. Again, we use Mistral-$7$B  to compute values for the $7$ features discussed in Section 4.1. Step-$2$: We ask humans (domain experts) for their preference, given reviews and two opinion summaries (A, B). We use the preference data and the features to train the reward model, $\varphi_{op}$. Step-$3$: We sample instances from PromptOpinSumm dataset; $\varphi_{op}$ assigns a score to the sampled summaries, the policy, $\pi_{\theta}$, assigns log probabilities to these summaries. Proximal Policy Optimization uses these to update $\pi_{\theta}$.
4 Efficient Reward Modeling
We highlighted in Section 1 how the reward model ($\varphi$) can depend on the downstream task. Such dependence necessitates task/domain-specific human preference datasets, which are costly and time-consuming to create. This creates an obstacle in employing RLHF in task/domain-specific setups, thus hindering the steering of LLMs towards task/domain-specific human values.
We solve this challenge by leveraging domain knowledge. The key insight is that we can use the domain knowledge to impart some inductive biases into the mathematical modeling of $\mathbf{\varphi}$. This would significantly reduce the amount of data required for training $\varphi$. Specifically, we say that $\varphi_{\tau}$ (reward model for a domain $\tau$) can be modelled by some numeric features $v_{1},v_{2},\cdots,v_{n}$. These $n$ features fully characterize [^7] Example of such characterization: Features like fluency, coherence, etc. can characterize text generated by an LLM. the outputs from the LLM on some input. Thus, instead of training $\varphi_{\tau}$ on the text data ($\{(X,Y_{w},Y_{l})\>|\>Y_{w}\succ Y_{l}\}$), we use the $n$ features. Such a formulation for $\varphi$ also brings interpretability and frees $\varphi$ from Alignment Tax.
In Section 4.1, we detail our technique for the task/domain of E-Commerce Opinion Summarization—the task of summarizing user reviews for a product. Typically, user reviews discuss several aspects of a product and opinions/sentiments towards these aspects. An opinion summary must reflect all the aspects discussed by the input reviews and the opinions expressed towards these aspects. We discuss how we leverage such desirable properties to model $\varphi$.
4.1 Inducing Domain Knowledge
We identify desirable properties in an opinion summary with the help of domain experts [^8] Domain experts are from an E-Commerce platform.. We held multiple discussions to finalize the set of desirable properties. We show that these properties are correlated to humans’ judgement of summary in Appendix A (Table 3). Based on these properties, we model $\varphi_{op}$ (reward model for opinion summarization) as: $\varphi_{op}$$\>=f(v)$, where $v\in$aspect-coverage, opinion-faithfulness, opinion-coverage, conciseness, relevance, hallucination, language-correctness}. The features aspect-coverage, opinion-faithfulness  and opinion-coverage  check if the generated opinion summary covers all mentioned aspects and opinions faithfully. The features conciseness, relevance , and hallucination  check if the generated summary is concise, relevant to the input reviews, and is free from hallucination. The feature language-correctness  checks if the generated text follows the language rules. We provide more details in Appendix A. These features, together, characterize the goodness of an opinion summary. We instruct Mistral-$7$B (Appendix A) to generate values for these features for an opinion summary, given reviews. We denote this transformation (from reviews and summary to $7$ features) using $\Phi$.
We train $\varphi_{op}$ using OpinPref, which is of the form: $\mathcal{D}_{h}=\{(R,s_{w},s_{l})\>|\>s_{w}\succ s_{l}\}$, where $R$ is the set of reviews and $s_{w}$ and $s_{l}$ are opinion summaries. We parameterize $\varphi_{op}$ using a Feed-Forward Neural Network and train it using the Elo-loss Ouyang et al. (2022); Glaese et al. (2022) (Equation 4.1; $\Phi(R,s_{i})$ uses Mistral-$7$B to compute the $7$ features; only $\varphi_{op}$ is trainable, $\Phi$ is not).
After such an efficient reward modeling, we use $\varphi_{op}$ for regular RLHF training (Appendix B) to get an Opinion Summarizer aligned with human goals. We illustrate the whole flow in Figure 2.
$\mathcal{L}_{pr}=-\mathbb{E}_{(R,s_{w},s_{l})\sim\mathcal{D}_{h}}\Big{[}$ $\log\sigma\big{(}\varphi_{op}(\Phi(R,s_{l}))$
$-\varphi_{op}(\Phi(R,s_{w}))\big{)}\Big{]}$ (1)
5 Experiments
We test our technique against the State-of-the-Art (SOTA) models, and strong Reinforcement Learning (RL) and RLHF baselines (our design and contemporary works). We list the questions we attempt to answer (through the experiments) in Section 5.1. We conduct automatic, human, and Gpt-$4$ evaluations to verify our claim. We find that our proposed technique excels significantly. In the rest of the section, we describe our models (Section 5.1) and evaluation results (Section 5.2).
5.1 Models & Objectives
We train the following models:
Supervised: This is a supervised model trained using Maximum Likelihood Estimation.
NaiveMean: This is a Reinforcement Learning model, where the reward is computed by averaging the feature values obtained using $\Phi$.
[…]
Inductive-Bias: This RLHF model is trained following our hypothesis (infusing domain knowledge into $\varphi$ to reap benefits of RLHF with modest human preference data). We train $\varphi_{op}$ using OpinPref dataset.
With these models, we ask the following questions in our experiments:
Scene-I: How effective is our technique (infusing domain knowledge into $\varphi$ to reap benefits of RLHF with modest human preference data) over and above the usage of a good training dataset? A comparative evaluation of Supervised and Inductive-Bias would answer this.
Scene-II: How effective is our technique over and above vanilla RL? A comparative evaluation of NaiveMean and Inductive-Bias would answer this.
Scene-III: How effective is our technique over contemporary RLHF techniques, which work without preference data? A comparative evaluation of Synth-Feedback and Inductive-Bias would answer this.
Scene-IV: How effective is our technique, agnostic of the preference data? This question is raised to answer whether the gains are solely due to the good quality of OpinPref, or the approach. A comparative evaluation between DPO (Rafailov et al. (2023), which uses OpinPref in a supervised fashion) and Inductive-Bias would answer this.
In addition to the above questions, we also check how our models fare against the SOTA (Op-Sum-Gen: Siledar et al. (2023a), Medos: Siledar et al. (2024), etc.). We do not use vanilla RLHF Ziegler et al. (2019); Bai et al. (2022a) as a baseline, as it requires huge human preference data. Given that the goal of the paper is not to propose a new RLHF technique, but rather to propose a way to use RLHF with modest human preference annotations, omitting vanilla RLHF as a baseline does not affect our conclusions in any way.
We use BART-Large Lewis et al. (2020) for all of our models. The choice of the model is governed by two factors: (a) It provides a similar environment (model size) for comparison with SOTA, (b) We find that LLMs (Mistral-$7$B, LLaMA2222-$7$B, Zephyr-$7$B, etc.) are already quite good at opinion summarization; thus any performance benefits (over SOTA) in those models cannot be reliably attributed to our approach. We include implementation details in Appendix E.
Model-Code Amazon Amazon-R Amazon-Rdq
R-$1\uparrow$ R-$2\>\uparrow$ R-L\>$\uparrow$ R-$1\>\uparrow$ R-$2\>\uparrow$ R-L\>$\uparrow$ R-$1\>\uparrow$ R-$2\>\uparrow$ R-L\>$\uparrow$
Prior Works MeanSum Chu and Liu (2018) $29.20$ $4.70$ $18.15$- -- -- -- -- --
CopyCat Bražinskas et al. (2020) $31.97$ $5.81$ $20.16$ $20.09$ $1.79$ $12.94$ $20.54$ $1.94$ $13.85$
PlanSum Amplayo and Lapata (2020) $32.87$ $6.12$ $19.05$ $20.49$ $1.76$ $12.44$ $19.09$ $1.58$ $12.02$
MultimodalSum Im et al. (2021) $34.19$ $7.05$ $20.81$ $21.43$ $1.58$ $13.20$ $20.39$ $2.08$ $12.83$
Op-Sum-GenSiledar et al. (2023a) $\mathbf{35.46}$ $\underline{7.30}$ $\mathbf{21.50}$- -- -- -- -- --
MedosSiledar et al. (2024) $\underline{34.63}$ $\mathbf{7.48}$ $\underline{20.97}$ $23.92$ $2.27$ $14.69$ $25.44$ $4.16$ $16.45$
Ours’ DPO $23.96$ $4.54$ $14.27$ $26.37$ $4.25$ $15.03$ $25.13$ $3.84$ $14.86$
Supervised $28.99$ $4.90$ $16.91$ $32.52$ $5.96$ $18.07$ $30.46$ $\underline{5.49}$ $17.63$
NaiveMean $28.08$ $4.81$ $16.77$ $\mathbf{34.0}$ $\underline{6.30}$ $\underline{18.81}$ $\underline{30.97}$ $5.25$ $\underline{18.36}$
Synth-Feedback $29.39$ $4.68$ $17.35$ $33.62$ $6.06$ $18.61$ $30.65$ $5.23$ $18.11$
Inductive-Bias $28.41$ $4.65$ $16.90$ $\underline{33.95}$ $\mathbf{6.40}$ $\mathbf{19.23}$ $\mathbf{31.89}$ $\mathbf{5.78}$ $\mathbf{18.84}$
Table 1: Reference-based Evaluation Results (R-$1$: Rouge-$1$, R-$2$: Rouge-$2$, R-L: Rouge-$L$) for the Amazon, Amazon-R and Amazon-Rdq benchmarks. We see the following things: (a) Our proposed dataset (PromptOpinSumm) leads to marked increased over the SOTA (by $\sim 4$ R-L points), (b) Inductive-Biasproves to be the winner in all the four scenarios: Scene-I, Scene-II, Scene-III and Scene-IV (Section 5.1), proving the efficacy of our technique. We also see that for the Amazon benchmark, our models lag behind. However, this is expected, as we highlight in Section 3.2.
5.2 Evaluation Results
We test our approach on $\mathbf{9}$ benchmarks (Section 3.2). In the main manuscript, we report automatic evaluation results on Amazon-based benchmarks (Table 1), human evaluation on the Amazon benchmark, and Gpt-$4$ evaluations on Amazon, Flipkart and Oposum+ benchmarks. Liu et al. (2023) show that GPT-$4$ evaluations correlate well with human evaluations for summarization; hence, in the interest of time and monetary expense, we resort to Gpt-$4$ evaluation. We include automatic evaluation results for the rest of the benchmarks (Tables 4 and 5), and BertScore based evaluations (Table 6) for all the benchmarks in Appendix C. We also include model generations for a randomly sampled product in Appendix C (Table 14). Due to the shortcomings highlighted in Section 3.2 and Appendix D, we complement our automatic evaluations of Amazon, Flipkart and Oposum+ with human and Gpt-$4$ evaluations.
Automatic Evaluation.  From Table 1, we see that our proposed models are always better than the SOTA for Amazon-R and Amazon-Rdq. Supervised Fine Tuning on PromptOpinSumm (Supervised model) helps achieve significantly better Rouge scores. This highlights the efficacy of our proposed PromptOpinSumm dataset. From the automatic evaluations on Amazon-R and Amazon-Rdq, we see the following things:
Answer to Scene-I: We see that Inductive-Bias achieves gains over Supervised. This answers the question in Scene-I: Our technique is effective over and above using a good dataset.
Answer to Scene-II: We see that Inductive-Bias achieves gains over NaiveMean. This answers the question in Scene-II: Our technique is effective over vanilla RL.
Answer to Scene-III: We see that Inductive-Bias achieves gains over Synth-Feedback. This answers the question in Scene-III: Our technique is effective over the SOTA RLHF technique, which works without human preference data.
Answer to Scene-IV: We see that Inductive-Bias achieves gains over DPO. This verifies that gains of Inductive-Bias can be safely attributed to the approach (not just the quality of OpinPref).
Human/Gpt-$4$ Evaluation.  We conduct human evaluation (Figure 1) for the Amazon benchmark, using $3$ domain experts (details in Appendix I). We observe a Fleiss’ Kappa ($\kappa$) score of $56.25\%$ (moderate agreement; agreement is moderate when $40\%\leq\kappa<60\%$). We ask the experts to rank the summaries (anonymized and shuffled) given the reviews. Given the rankings, we compute the fraction of pairwise wins, ties, and losses among all the models. We compare summaries from Supervised, NaiveMean, Synth-Feedback, Inductive-Bias, Op-Sum-Gen (SOTA) models and ground truth summaries. We include ground truth summaries in the evaluation to verify our claims about the quality of the benchmarks. From Figure 1, we see that Inductive-Bias wins significantly over the competitors, further proving the efficacy of our technique.
We run Gpt-$4$ evaluations for Amazon, Flipkart and Oposum+ benchmarks (Figures 3, 4, 5). We run Gpt-$4$ evaluations for Amazon, as the agreement in human evaluation was moderate. We arrive at the same conclusions as human evaluation. We prompt Gpt-$4$ to rank the summaries (anonymized and shuffled) given the reviews. As before, we compute the fraction of wins, ties, and losses. Again, we see that Inductive-Bias remains a clear winner.
Figure 3: Gpt-$4$ Eval: Pairwise win-tie-loss percentage of Inductive-Bias model vs. competitors, for Amazon benchmark.
Figure 4: Gpt-$4$ Eval: Pairwise win-tie-loss percentage of Inductive-Bias model vs. competitors, for Flipkart benchmark. Note that for the Flipkart benchmark, we do not have results from Op-Sum-Gen, as Siledar et al. (2023a) only provide aspect-specific summaries.
Figure 5: Gpt-$4$ Eval: Pairwise win-tie-loss percentage of Inductive-Bias model vs. competitors, for Oposum+ benchmark.
6 Analysis
We perform a two-fold analysis: (a) First, we see the domain knowledge features influence for $\varphi_{op}$, (b) Second, we see how the ground truth summary and summary from trained models fare on the domain knowledge features. This two-fold analysis helps us understand: (a) which features influence the latent reward model within humans [^9] Note that the trained $\varphi_{op}$ represents latent human reward model. the most, and (b) how the ground truth summary and summary from trained models fare on these influential features. Performing well on influential features would mean the summary aligns well with the latent reward model within humans.
6.1 Analysis of $\varphi_{op}$
$\varphi_{op}$ model has been trained on a set of features specified by domain experts. We analyze the relative influence of each feature on the final score assigned by $\varphi_{op}$. Doing this helps us understand an approximate importance [^10] We call this approximate importance as the influence of a feature on the output is not necessarily its importance. of each of these features. We do this by varying each feature by $\delta\>(=0.1)$ while keeping the other features constant, over multiple possible values of all features (Equation 6.1).
$\Delta_{i}=\dfrac{1}{2\delta}$ $\sum_{\mathbf{x}}\big{(}f(x_{1},\cdots,x_{i}+\delta,\cdots,x_{n})$
$-f(x_{1},\cdots,x_{i}-\delta,\cdots,x_{n})\big{)}$ (2)
Figure 6 highlights the features’ relative influence. We see that hallucination is most influential. This aligns with what our human preference annotators report—hallucination in summary is the primary cause of rejection. We see that hallucinations are mostly within the opinions in the summary. This is also reflected in Figure 6: opinion-faithfulness has significant influence. We also see that annotators prefer summaries with more specifics, i.e. they include more aspects: aspect-coverage has significant influence.
6.2 Analysis of Summaries
We analyze the top-$3$ performing models (in human and Gpt-$4$ evaluations) for the following features: opinion-coverage, opinion-faithfulness, hallucination and relevance. We show the analysis only for the Amazon benchmark in the main manuscript, we include the rest in Appendix J. Table 2 shows the performance on these features. We see that Inductive-Bias model fares much better than the competitors on hallucination (the most influential metric). For relevance, aspect-coverage and opinion-faithfulness, our model is fairly better than the other models.
Figure 6: Relative Influence of all features in $\varphi_{op}$. All the influences sum to $1$.
Models AC $\uparrow$ OPF $\uparrow$ RE $\uparrow$ HL $\uparrow$
IB $\mathbf{3.60}$ $\mathbf{3.93}$ $\mathbf{4.06}$ $\mathbf{4.07}$
SF $3.43$ $3.73$ $4.04$ $3.94$
NM $3.57$ $3.91$ $4.04$ $3.09$
Table 2: Scores on domain knowledge-based features (AC: aspect-coverage, RE: relevance, OPF: opinion-faithfulness, HL: hallucination) on the Amazon benchmark for top-$3$ models (IB: Inductive-Bias, NM: NaiveMean, SF: Synth-Feedback). Note that for hallucination, $\Phi$ gives a higher score for less hallucination in the text.
This shows that our technique helps Inductive-Bias model perform well on features that influence the latent reward model within humans for opinion summarization. This means that our technique helps Inductive-Bias model achieve a significant alignment with the latent reward model. This conclusion verifies our hypothesis (in the domain of opinion summarization): A domain-knowledge infused reward model ($\varphi_{op}$) can help achieve alignment with latent reward model of humans for a task, with modest human preference annotations.
7 Summary, Conclusion and Future Work
In this work, we propose a novel Reward Modeling technique via Domain Knowledge Infusion. We verify our approach for E-Commerce Opinion Summarization, where we achieve State-of-the-Art, while significantly reducing the amount of human preference annotations required (just $940$ samples). In addition to advancing SOTA and reducing preference annotations, our technique provides another two-fold benefits: (a) No Alignment Tax and (b) Interpretability. Due to the interpretable nature, we find that our model does achieve alignment with human goals for Opinion Summarization through analysis. From the results and analysis, we conclude that Domain Knowledge Infusion into Reward Modeling is a viable solution to reduce human preference annotations for downstream tasks. In the future, we will verify this for other domains.
8 Ethical Considerations
We contribute two datasets in our work: PromptOpinSumm, OpinPref. These datasets are generated using an open-source model Mistral-$7$B Jiang et al. (2023). We would release the datasets to further research in Opinion Summarization. For the OpinPref, to the best of our knowledge, we have seen that it does not contain any harmful content, such as social biases, stereotypes, etc. However, we have seen that it contains products of explicit nature (sexual products). For the PromptOpinSumm dataset, to the best of our knowledge, there is no presence of harmful content, such as social biases, stereotypes etc. We urge the research community to use the datasets with caution and check for potential harmfulness, based on their use-cases.
9 Limitations
A limitation of our work is we have tested our approach for one domain: Opinion Summarization. However, we do not believe that this weakens our argument, as we have exhaustively shown that our approach not only advances SOTA but also interpretably achieves alignment with humans. Future work in other domains would help in verifying this claim for other domains. Another limitation is: we see empirically that $\Phi$ works well for Opinion Summarization, to extract the scores for the $7$ features. However, there is no guarantee that such out-of-the-box performance would be reflected in another domain. Some fine-tuning might be necessary.
Appendix A Features for Reward Modeling
We use $7$ domain specific features for the reward model $\varphi_{op}$. We identify these features after extensive discussions with the domain experts. For each feature we prompt Mistral-$7$B to generate a score within 00 and $5$. We give elongated instructions, including rough rubriks, and rules to generate the scores. This is a reason why we use an instruction-tuned model. For each feature, 00 means the model is doing bad on the feature, and $5$ means the model is doing good on the feature. We define all the features below:
aspect-coverage: This feature considers the aspect coverage within an opinion summary. The feature assumes a value $5$ if all the aspects of the product, mentioned in the reviews, are mentioned in the summary. If none of the aspects are picked, the feature assumes a value 00.
opinion-faithfulness: This feature considers whether the mentioned opinions/sentiments in the summary are correct, that is, they are picked correctly from the reviews. For example, if an user mentions that they are happy with the battery of a phone, and the summary mentions that users are unhappy with the battery, the summary will not be considered faithful to opinion in the review. The feature assumes a value $5$ if all the opinions are faithfully reflected. If no opinion is faithfully reflected, the value would be 00.
opinion-coverage: This feature considers whether all the opinions in the input reviews are picked by the opinion summary. The feature assumes a value $5$ if all the opinions are picked up. If none of the opinions are picked up, the feature assumes a value 00.
relevance: This feature checks if the summary is relevant to the input reviews (that is the product). The feature assumes a value $5$ if summary is completely relevant. If it is completely irrelevant, the feature assumes a value 00.
conciseness: This feature considers the conciseness and completeness of the opinion summary. The feature assumes a value $5$ if the summary is concise and complete—not one phrase/sentence can be dropped off. It assumes a value 00 if the summary is totally incomplete, or very verbose.
hallucination: This feature considers the factuality of the opinion summary. The feature assumes a value $5$ if the summary is totally factual, with respect to the input reviews. If there are a lot of hallucinations, the feature assumes a value 00.
language-correctness: This feature checks the correctness of language/text in the opinion summary. The feature assumes a value $5$ if the summary is grammatically fully correct. It assumes a value 00 if the summary is very poor linguistically.
For conciseness, we do not include the prompts in the paper, we would release them as separate artifacts, with the datasets, in the camera ready version.
We also analyze how these features correlate with humans’ judgement of goodness of opinion summaries. We do this by looking at the scores for these features for preferred and dis-preferred summaries in the OpinPref dataset. In Table 3, we see that the preferred summaries clearly have a higher score on all the features, than the dis-preferred ones. This shows that the scores correlate well with humans’ judgement of goodness.
Feature Pref. Dis-pref.
aspect-coverage ($\uparrow$) $3.69$ $2.84$
opinion-faithfulness ($\uparrow$) $4.02$ $3.05$
opinion-coverage ($\uparrow$) $3.92$ $3.22$
conciseness ($\uparrow$) $4.05$ $3.44$
relevance ($\uparrow$) $4.10$ $3.10$
hallucination ($\uparrow$) $3.99$ $2.79$
language-correctness ($\uparrow$) $4.50$ $3.32$
Table 3: Scores for the domain knowledge based features. We see that for all the features, the human preferred (Pref.) summaries have higher scores than the ones rejected by humans (Dis-pref.). This shows that these features correlate well with humans’ judgement of goodness of an opinion summary.
Model-Code Flipkart Flipkart-R Flipkart-Rdq
R-$1\uparrow$ R-$2\>\uparrow$ R-L\>$\uparrow$ R-$1\>\uparrow$ R-$2\>\uparrow$ R-L\>$\uparrow$ R-$1\>\uparrow$ R-$2\>\uparrow$ R-L\>$\uparrow$
MedosSiledar et al. (2024) $25.97$ $\mathbf{5.29}$ $\underline{16.05}$ $26.29$ $4.03$ $16.59$ $22.92$ $4.30$ $16.35$
Ours’ DPO $28.85$ $4.10$ $15.55$ $34.23$ $7.86$ $18.62$ $29.96$ $5.25$ $17.28$
Supervised $27.38$ $4.09$ $15.37$ $39.32$ $10.52$ $22.56$ $32.25$ $6.88$ $19.04$
NaiveMean $28.34$ $\underline{4.38}$ $\mathbf{16.20}$ $\mathbf{40.56}$ $10.68$ $22.74$ $\underline{32.57}$ $6.67$ $\underline{19.39}$
Synth-Feedback $26.37$ $4.18$ $15.48$ $38.77$ $\underline{10.99}$ $\underline{22.97}$ $31.04$ $\underline{6.98}$ $18.59$
Inductive-Bias $\mathbf{27.42}$ $4.21$ $15.71$ $\underline{39.10}$ $\mathbf{11.03}$ $\mathbf{23.30}$ $\mathbf{33.08}$ $\mathbf{7.30}$ $\mathbf{19.46}$
Table 4: Reference-based Evaluation Results (R-$1$: Rouge-$1$, R-$2$: Rouge-$2$, R-L: Rouge-$L$) for the Flipkart, Flipkart-R and Flipkart-Rdq benchmarks. We see the following things: (a) Our proposed dataset (PromptOpinSumm) leads to marked increased over the SOTA (Medos; by $\sim 6$ R-L points), (b) Inductive-Biasproves to be the winner in all the four scenarios: Scene-I, Scene-II, Scene-III and Scene-IV (Section 5.1), proving the efficacy of our technique. We also see that for Flipkart benchmark, despite the shortcomings, our models perform similar to the SOTA.
Model-Code Oposum+ Oposum-R Oposum-Rdq
R-$1\uparrow$ R-$2\>\uparrow$ R-L\>$\uparrow$ R-$1\>\uparrow$ R-$2\>\uparrow$ R-L\>$\uparrow$ R-$1\>\uparrow$ R-$2\>\uparrow$ R-L\>$\uparrow$
Prior Works MeanSum Chu and Liu (2018) $26.25$ $4.62$ $16.49$- -- -- -- -- --
CopyCat Bražinskas et al. (2020) $27.98$ $5.79$ $17.07$ $22.41$ $2.30$ $13.94$ $22.38$ $2.03$ $14.06$
PlanSum Amplayo and Lapata (2020) $30.26$ $5.29$ $17.48$ $22.37$ $2.05$ $13.32$ $22.64$ $2.25$ $13.71$
MultimodalSum Im et al. (2021) $33.08$ $7.46$ $19.75$ $23.35$ $2.98$ $14.53$ $23.73$ $2.80$ $14.70$
Op-Sum-GenSiledar et al. (2023a) $36.44$ $8.50$ $22.03$ $25.65$ $3.56$ $15.83$ $24.66$ $3.25$ $15.54$
MedosSiledar et al. (2024) $\mathbf{36.57}$ $\underline{8.79}$ $\mathbf{21.35}$ $26.82$ $3.67$ $15.92$ $26.32$ $3.34$ $16.10$
Ours’ DPO $27.64$ $7.34$ $16.50$ $33.69$ $6.62$ $18.55$ $30.95$ $5.89$ $17.60$
Supervised $30.57$ $8.02$ $16.90$ $38.32$ $9.10$ $20.35$ $35.69$ $8.17$ $19.28$
NaiveMean $31.47$ $8.0$ $16.99$ $40.16$ $9.84$ $21.74$ $35.90$ $8.33$ $20.13$
Synth-Feedback $\underline{31.66}$ $\mathbf{8.86}$ $\underline{17.91}$ $\underline{41.32}$ $\mathbf{10.40}$ $\mathbf{22.23}$ $\mathbf{37.85}$ $\underline{8.94}$ $\underline{20.71}$
Inductive-Bias $31.15$ $8.15$ $17.46$ $\mathbf{41.58}$ $\underline{10.32}$ $\underline{22.02}$ $\underline{37.56}$ $\mathbf{9.21}$ $\mathbf{20.88}$
Table 5: Reference-based Evaluation Results (R-$1$: Rouge-$1$, R-$2$: Rouge-$2$, R-L: Rouge-$L$) for the Oposum+, Oposum-R and Oposum-Rdq benchmarks. We see the following things: (a) Our proposed dataset (PromptOpinSumm) leads to marked increased over the SOTA (Medos; by $\sim 6$ R-L points), (b) Inductive-Biasproves to be the winner in almost all of the four scenarios: Scene-I, Scene-II, Scene-III and Scene-IV (Section 5.1), proving the efficacy of our technique. We also see that for Oposum+ benchmark, our models lag behind. However, this is expected, as we highlight in Section 3.2.
Model Code Amazon Amazon-R Amazon-Rdq Oposum+ Oposum-R Oposum-Rdq Flipkart Flipkart-R Flipkart-Rdq
Op-Sum-Gen $\mathbf{88.78}$ $86.94$ $86.76$ $\mathbf{86.63}$ $86.96$ $86.95$- -- --
Siledar et al. (2023a)
DPO $86.45$ $86.60^{\dagger}$ $86.37^{\dagger}$ $84.39$ $87.35^{*}$ $86.90$ $83.75$ $86.61$ $85.40$
Supervised $87.79$ $88.23^{*}$ $87.76^{*}$ $85.13$ $88.59^{*}$ $88.02^{*}$ $84.21$ $88.11$ $86.40$
NaiveMean $87.95$ $\underline{88.29}^{*}$ $\underline{87.81}^{*}$ $85.25$ $88.96^{*}$ $88.39^{*}$ $\underline{84.32}$ $\underline{88.29}$ $\underline{86.52}$
Synth-Feedback $87.81$ $88.28^{*}$ $87.74^{*}$ $85.22$ $\underline{89.08}^{*}$ $\underline{88.45}^{*}$ $84.27$ $88.28$ $86.49$
Inductive-Bias $\underline{87.98}$ $\mathbf{88.41}^{*}$ $\mathbf{88.16}^{*}$ $\underline{85.33}$ $\mathbf{89.09}^{*}$ $\mathbf{88.46}^{*}$ $\mathbf{84.33}$ $\mathbf{88.34}$ $\mathbf{86.61}$
Table 6: BertScore evaluation results on the $9$ benchmark datasets. We observe a similar trend as Rouge evaluations: SOTA is better than our models for the Amazon and Oposum+ benchmarks, which is expected (Section 3.2). For the rest of the datasets, we see that our models are significantly better. We do not include SOTA results for Flipkart-based benchmarks, as Op-Sum-Gen only provide aspect-specific summaries for the same. denotes gain is statistically significant compared to SOTA with significance level $1\%$, denotes gain is statistically significant compared to SOTA with significance level $5\%$.
Appendix B RLHF Training Pipeline
Using the trained reward model, we follow a similar training pipeline as Bai et al. (2022a); Ouyang et al. (2022), with a modification: Limited Trajectory Reinforcement Learning. Computing the transformation $\Phi$ for each generation online (during training) is expensive, especially with limited compute resources. To circumvent this, we limit the trajectories that are explored by our policy, $\pi_{\theta}$. Specifically, we limit it to the Good-Sum, SBad-Sum and VBad-Sum trajectories in the PromptOpinSumm dataset. Having varying levels of quality in PromptOpinSumm is of use here—it lets the model still explore trajectories of several quality. Thus, we have an offline experience buffer, with $\Phi$ precomputed, for $\pi_{\theta}$ learn from.
We use Proximal Policy Optimization (PPO) Schulman et al. (2017) to train our model (Equation B). For each training step, we sample ($R,s,\Phi(R,s)$) tuples from PromptOpinSumm. We use the trained $\varphi_{op}$ to compute the reward for $s$ ($=\varphi_{op}\big{(}\Phi(R,s)\big{)}$). PPO uses this to update the log probability assigned by $\pi_{\theta}$. We parameterize $\pi_{\theta}$ using a Transformer model, which takes reviews as input, and generates an opinion summary.
$\mathcal{L}_{PPO}=-\mathbb{E}$ [${}_{(R,s,\Phi(s))}\bigg{[}\varphi_{op}(\Phi(R,s))$
$-\beta\log\Big{(}\dfrac{\pi^{RL}_{\theta}(s|R)}{\pi^{SFT}(s|R)}\Big{)}\bigg{]}$ (3)
Appendix C Additional Automatic Evaluation Results
In addition to the Amazon-based benchmarks (Table 1), we also report results for Flipkart and Oposum+ based benchmarks (Tables 4 and 5). As before, we see that Inductive-Bias is almost always the winner. As before, we draw similar conclusions for Scene-I, Scene-II and Scene-III: Inductive-Bias wins, further strengthening the conclusion that our methodology is effective. We also see that, inspite of the shortcomings of the Flipkart benchmark, our models perform similar to the SOTA.
We also include BertScore evaluations for all the $9$ benchmarks in Table 6. We see similar trends as Rouge Evaluation: our models are significantly better than the SOTA in majority of the benchmarks.
For a qualitative understanding, we include generations from several models on a randomly picked sample from the Amazon benchmark in Table 14.
Appendix D Details on the Benchmark Datasets
In this section we discuss details about the benchmarks, such as the domain of the products, summary statistics and finally highlight some shortcomings in the Amazon, Oposum+ and Flipkart datasets. Amazon has reviews for $32$ products from $4$ domains: “electronics”, “home & kitchen”, “personal care”, and “clothing, shoes & jewellery”. Oposum+ has reviews for $60$ products from $6$ domains: “laptop bags”, “bluetooth headsets”, “boots”, “keyboards”, “television”, and “vacuums”. Flipkart has reviews for $147$ products from $3$ domains: “laptops”, “mobiles”, and “tablets”. Table 7 includes summary statistics for the benchmarks.
Characteristic Oposum+ Amazon Flipkart
# domains $6$ $4$ $3$
# products $60$ $32$ $147$
# reviews $10$ $8$ $10$
per product
# summaries $3$ $3$ $1$
per product
Table 7: Statistics of the benchmark datasets. Oposum+ represents the statistics of all Oposum+ based benchmarks (Oposum+, Oposum-R and Oposum-Rdq). Similar is the case for Amazon and Flipkart.
Finally, now we highlight the shortcomings of the benchmark datasets in the rest of the discussion.
Amazon: Bražinskas et al. (2020) designed the test-set in such a way that the summary has to read like a review, for instance, summary would contain ‘I think the quality has come down over the years.’, instead of ‘Users think that quality has come down over years’. Due to this writing style, the summaries read like reviews and are often in first person—high overlap would not necessarily mean a better summary, it would rather mean a better review.
Flipkart: Siledar et al. (2023b) generate this dataset by listing out the aspect-wise pros and cons presented within the reviews. We form an opinion summary by concatenating these pros and cons. Due to this, the summaries have frequent incoherent sentences.
Oposum+: Amplayo et al. (2021) create this benchmark by extracting sentences from the input reviews. Hence, this dataset has similar drawbacks as the Amazon benchmark.
Amazon
    • Nice boots but run a bit narrow. They look great but I think the quality has come down over the years. Still comfortable but I wish they broke in easier. I recommend these for any lady who is patient and looking for comfort.
Oposum+
    • great product for the cost . very easy to use and compatible with all of my phones ! it holds a charge great , is light enough and fits perfectly in my ear . the sound quality is great , the style is very cool and the unit feels top quality . it would drop and reconnect every 10 seconds nobody could hear me i could n’t get it to unpair from the phone , there ’s apparently no noise-cancellation in these . the battery life is … bizarre . cheap , plastic-y , and poor sound quality .
Flipkart
  • Summary
    • Pros
      • Design: The full-metal Infinix INBook X1 Core i3 has a top notch and premium design.
      • 35.56 cm (14 inch) 1920 x 1080 Pixel Full HD IPS Display: 100% sRGB with 300nits brightness ensures an excellent display.
      • Battery: Long-lasting battery. Gives around 8 hours of backup on normal usage.
      • Performance: The combination of Intel Core processor chip, high RAM size and sufficient storage capacity gives this laptop a high-speed performance experience.
      • Price: "Totally worth it in this price range.
    • Cons
      • Charging: Some current leakage during charging. Sometimes the laptop won’t charge.
      • Trackpad: Not upto the mark.
    • Verdict: This laptop comes with a i3 10th gen dual core processor which is suitable for normal tasks like web browsing, online classes and watching movies. Not recommended as a gaming laptop.
    • Additional Information: Can handle video editing and expandable SSD.
Appendix E Implementation Details
We use BART-Large Lewis et al. (2020) as our policy ($\pi_{\theta}$) in all of the models. We do this to have a fair comparison with the state-of-the-art in Opinion Summarization. We use AdamW Optimizer to train the models, with a weight decay of $0.05$. We use a cosine learning rate scheduler. We run a hyperparameter sweep on batch size, learning rate, and learning rate warmup. We include the possible values for the sweep in Table 8. We train all of our models using 2×2$\times$2 × A100100100100 GPUs ($80$GB)
Hyperparameter Values
batch size [$64$, $128$, $256$]
learning rate $\sim\mathcal{U}(5e^{-6},5e^{-5})$
learning rate warmup $\sim\mathcal{U}(0.2,0.4)$
Table 8: Possible Values for Hyperparameters. For learning rate warmup, we sample the fraction of total steps the learning should be warmed up. For example, if the learning rate warmup is $0.2$, it means that the learning rate will have a linear warmup for $20\%$ of the total training steps.
For the reward model, $\varphi_{op}$, we use a Feed Forward Network for the Policy Model. We use AdamW Optimizer to train the models, with a weight decay of $0.05$. As before, we run a hyperparameter sweep on batch size, learning rate, and learning rate warmup. Table 9 includes details on the hyperparameters.
Hyperparameter Values
batch size [$32$, $64$, $128$]
learning rate $\sim\mathcal{U}(5e^{-3},1e^{-1})$
Table 9: Possible Values for Hyperparameters for the Reward Model. For learning rate warmup, we sample the fraction of total steps the learning should be warmed up. For example, if the learning rate warmup is $0.2$, it means that the learning rate will have a linear warmup for $20\%$ of the total training steps.
Appendix F Generated Summary Lengths
We analyze the generation lengths of the models, and the ground truth summary. Table 10 lists the summary lengths. We see that theDPO model generates very verbose summary. Additionally, we also see that the Inductive-Bias model generates very concise summaries.
Model Amazon Oposum+ Flipkart
Ground-Truth $60.65$ $85.86$ $129.91$
NaiveMean $91.09$ $114.67$ $75.48$
Synth-Feedback $80.31$ $115.37$ $71.11$
Op-Sum-Gen $55.84$ $62.93$
Inductive-Bias $81.62$ $88.63$ $73.57$
Supervised $81.31$ $117.03$ $74.56$
DPO $138.50$ $141.50$ $131.40$
Table 10: Generation Length Statistics: number of words in summaries. We use Nltk to tokenize the text.
Good: The users have shared their positive experiences with the protective covers for Macbooks. They appreciate the ease of application, vibrant colors, quick shipping, and great quality. Some have mentioned the covers fit their Macbooks perfectly, while others have shared their disappointment when they realized it did not fit their specific model. A few users have noted the covers add a personal touch and pizzaz to their devices. However, some users have reported issues with the covers not fitting their Macbooks or falling off, leaving them feeling disappointed and frustrated.
Slightly bad: Some users have reported positive experiences with the Macbook cover, praising its protective qualities, easy application, and vibrant colors. Others, however, have encountered issues with it not fitting properly on their devices or being returned due to size mismatches. Overall, the product has been described as cute, great quality, and worth the investment, though some buyers have experienced disappointment with its size compatibility and easy detachment.
Very Bad: This is a terrible product for people with old Macbooks, it doesn’t fit at all! And don’t even bother looking at the description carefully before buying.
Table 11: Example summaries from PromptOpinSumm.
Appendix G Details on PromptOpinSumm
Here we provide more details on the generated PromptOpinSumm dataset. Table 13 includes summary statistics of the generated dataset. We include an example from PromptOpinSumm dataset (Table 11). We show one sample from Good-Sum, SBad-Sum and VBad-Sum quality each. We do not include the reviews for conciseness. However, we incluce the salient aspects of the reviews. The reviews talk about the following things:
  1. 1.
    Great price, Nice looking / Good color, Good utility / Good protection, Quick shipping, Nice fitting, Good accessibility of the laptop while the cover is on, Good finish quality.
  2. 2.
    Not a good fit for older macbooks, Broken / Unusable for the original ($\sim 2009$ $2010$) white macbooks, cannot return return.
Good-Sum manages to discuss all of these things—it was able to detect that the cover does not fit specific models (highlighted in blue). It also detects that a few users like the “personal touch addition” factor of the cover.
VBad-Sum is also a good representation of a bad summary—it totally ignores the “positive” aspects of the product and presents the “negative” aspects only. It fails at Aspect Coverage, Opinion Faithfulness and Opinion Coverage.
SBad-Sum maintains almost a similar quality as the Good one. However, it fails to draw out certain aspects, such as “pizzaz”, “personal touch addition”, etc.
Appendix H Statistics of the OpinPref dataset
We look at the summary statistics for the OpinPref dataset. Table 12. We see that, interestingly, annotators prefer longer summaries—this is because these summaries contain more specifics and details from the reviews.
Characteristic Value
# words in reviews $641.21$
# reviews $13.08$
# words in summaries $73.16$
# words in preferred summaries $85.41$
# words in unpreferred summaries $66.91$
Table 12: Statistics of the OpinPref dataset. We use Nltk to tokenize the text.
Split Characteristic $\mu$ $\sigma$
train # reviews $13.24$ $10.07$
# summaries $8.90$ $0.34$
# words in review $49.0$ $10.78$
# words in summary $78.28$ $34.45$
validation # reviews $10.53$ $6.80$
# summaries $8.98$ $0.16$
# words in review $48.65$ $10.63$
# words in summary $74.26$ $34.27$
Table 13: Statistics of PromptOpinSumm dataset. We use Nltk to tokenize the text.
Reviews: i really like these boots. they can be a beast to get on, like any boot fit to last; once on, they are incredibly comfortable. I have had them for 4 years and they still look great - worn in but not worn out… What great boots! They do take some breaking in - like all real leather does. Run maybe a tad small in the general spectrum, but hey, remember the break in period. Look fabulous, will last me for the next twenty years…. They bridge the fashion vs. function quandary beautifully. Every woman needs a pair of these Frye boots! Fit is true to normal size. I wear a standard 9. These fit in the shaft for any ladies who have bigger calves as I do. You have to fight to get them on first few times and they wear stiff until fully broken in. Worth the money and break in time. Boots excellent quality, beautiful skin, a modern model. But! they are very narrow! This is the second pair of shoes is a brand and my 8.5 size fits me perfectly, even he little more than usual. But these are very narrow! go to a cobbler, let him try to stretch. Be careful! bought these for the wife to have something to kick around in and keep up with the kids, but she doesn’t like them. Her heel lifts with every step, and they’re stiff and heavy. Not an everyday casual shoe. Don’t get me wrong, they’re really good boots, but they used to make them better. My last pair from a few years ago were much higher quality. Thicker sole, better full grain leather, now they only make the black in ’grease’, which stains all my socks black. Keeping them, but still disappointed in Frye. i bought it just for the heavy snow cuz my UGG got wet through.even though it is a little expensive, but the boots are nice and it was shipped to me for only two days.!!!!!! i love it They were comfortable the moment I put them on. The sizes are exact. The color the way they are made are what I expect from Frye which I’ve been wearing since the 70’s. With a 20year break from not wearing Frye it was a pleasant surprise the quality has stood the test of time.
Supervised: These Frye boots are a popular choice for women due to their comfort and durability. However, they can be a bit challenging to put on and require a break-in period. Some users have reported that they run small in size and are narrow, requiring a cobbler’s assistance for a proper fit. The leather quality is generally good, but some users have noted that the leather may not last as long as they had hoped.
NaiveMean: Users generally appreciate the comfort and durability of the Frye boots, with some mentioning the need for a break-in period. Sizes run true to normal, but some users find them to be a bit narrow. The boots are praised for their stylish appearance and ability to bridge the fashion and function quandary. However, some users have noted that they take some time to break in and are not suitable for everyday casual wear.
Synth-Feedback: These Frye boots are praised for their comfort and durability, with some users noting the need for a break-in period. However, some users have noted that they run small in size and require stretching, while others find them narrow and stiff. The leather is described as beautiful and of good quality, with a modern model. Some users have had them for several years and have not worn them out.
Inductive-Bias: These Frye boots are praised for their comfort and durability, with users expressing satisfaction with their fit and appearance. However, some users have noted that they require a break-in period and can be a bit narrow. The leather quality is praised for its durability and modern model, with some users expressing disappointment with the lack of improved quality in recent years.
Table 14: Example generation (randomly sampled) for some input reviews from all the models. Olive implies faithful/correct generation, while red indicates hallucinated text, or repetition. We see that only Inductive-Bias is free from red text. The model closest in performance to Inductive-Bias, the NaiveMean model, misses out on two aspects: leather-quality and quality-degradation. Inductive-Bias covers both, while being concise. We do not include DPO model in this comparison, as it was too verbose.
Appendix I Annotator Details
We include two disjoint sets of annotators in our work—first for creation of OpinPref ($3$ annotators), second for human evaluation  ($3$ annotators). For both annotations, we use domain experts. The domain experts are NLP researchers (age group: $24-30$) who have worked in Opinion Summarization for a long time, with publication experience (in A/A conferences). The domain experts for human evaluation also have a similar profile. The annotators have been paid generously, based on the standard annotation rates in the geographical location.
Appendix J All Evaluation Results
We include all of the evaluation results in this section. In Tables 15, 16, 17 and 17 we include pairwise comparison results, in a win/tie/loss format. We also include results on evaluation on how the models perform on the domain features in Tables 19, 20 and 21.
$\cdot$ Supervised NaiveMean Synth-Feedback Inductive-Bias Op-Sum-Gen
NaiveMean $0.50/0.06/0.38$
Synth-Feedback $0.44/0.12/0.44$ $0.40/0.09/0.5$
Inductive-Bias $\mathbf{0.56}/0.09/0.28$ $\mathbf{0.46}/0.18/0.31$ $\mathbf{0.56}/0.12/0.28$
Op-Sum-Gen $0.31/0.28/0.38$ $0.25/0.12/0.56$ $0.25/0.21/0.5$ $0.25/0.06/\mathbf{0.68}$
Ground-Truth $0.46/0.06/0.48$ $0.31/0.18/0.44$ $0.40/0.15/0.40$ $0.28/0.09/\mathbf{0.59}$ $0.5/0.09/0.38$
Table 15: Pairwise Win/Tie/Loss Results for all models in Human Evaluation for Amazon benchmark. We format the data as: win/tie/loss, win specifies how many time the row won over the column.
$\cdot$ Supervised NaiveMean Synth-Feedback Inductive-Bias Op-Sum-Gen
NaiveMean $0.63/0.12/0.25$
Synth-Feedback $0.59/0.12/0.28$ $0.5/0.06/0.44$
Inductive-Bias $\mathbf{0.62}/0.12/0.25$ $\mathbf{0.46}/0.09/0.44$ $\mathbf{0.5}/0.06/0.44$
Op-Sum-Gen $0.06/0.03/0.9$ 0.09/$0.0/0.9$ 0.0 / 0.90 $0.12/0.09/0.78$ $0.06/0.0/\mathbf{0.93}$
ground-truth $0.12/0.06/0.81$ $0.09/0.06/0.84$ $0.16/0.06/0.78$ $0.09/0.0/\mathbf{0.90}$ $0.68/0.09/0.22$
Table 16: Pairwise Win/Tie/Loss Results for all models in Gpt-$4$ Evaluation for Amazon benchmark. We format the data as: win/tie/loss, win specifies how many time the row won over the column.
$\cdot$ Supervised NaiveMean Synth-Feedback Inductive-Bias
NaiveMean $0.57/0.12/0.30$
Synth-Feedback $0.57/0.06/0.36$ $0.52/0.12/0.36$
Inductive-Bias $\mathbf{0.63}/0.12/0.25$ $\mathbf{0.54}/0.16/0.30$ $\mathbf{0.57}/0.08/0.34$
Ground-Truth $0.10/0.06/0.84$ $0.06/0.01/0.92$ $0.07/0.01/0.91$ $0.06/0.02/\mathbf{0.91}$
Table 17: Pairwise Win/Tie/Loss Results for all models in Gpt-$4$ Evaluation for Flipkart benchmark. We format the data as: win/tie/loss, win specifies how many time the row won over the column.
$\cdot$ Supervised NaiveMean Synth-Feedback Inductive-Bias Op-Sum-Gen
NaiveMean $0.56/0.03/0.4$
Synth-Feedback $0.5/0.16/0.34$ $0.46/0.1/0.44$
Inductive-Bias $\mathbf{0.66}/0.0/0.33$ $\mathbf{0.46}/0.1/0.44$ $\mathbf{0.56}/0.06/0.36$
Op-Sum-Gen $0.1/0.06/0.83$ $0.06/0.03/0.9$ $0.03/0.03/0.93$ $0.03/0.03/\mathbf{0.93}$
Ground-Truth $0.13/0.13/0.73$ $0.1/0.033/0.8666$ $0.06/0.06/0.86$ $0.06/0.06/\mathbf{0.86}$ $0.7/0.1/0.2$
Table 18: Pairwise Win/Tie/Loss Results for all models in Gpt-$4$ Evaluation for Oposum+ benchmark. We format the data as: win/tie/loss.
$\cdot$ AC OPF OPC CC RL HL LC
Supervised $3.43\pm 0.20$ $3.71\pm 0.37$ $3.67\pm 0.26$ $3.79\pm 0.31$ $4.04\pm 0.37$ $3.89\pm 0.39$ $4.55\pm 0.35$
NaiveMean $3.56\pm 0.22$ $3.91\pm 0.50$ $3.76\pm 0.38$ $3.89\pm 0.36$ $4.04\pm 0.48$ $3.99\pm 0.48$ $4.60\pm 0.27$
Synth-Feedback $3.55\pm 0.40$ $3.87\pm 0.71$ $3.71\pm 0.43$ $3.94\pm 0.50$ $4.04\pm 0.61$ $3.94\pm 0.68$ $4.38\pm 0.92$
Inductive-Bias $\mathbf{3.60}\pm\mathbf{0.17}$ $\mathbf{3.95}\pm\mathbf{0.40}$ $\mathbf{3.85}\pm\mathbf{0.25}$ $3.99\pm 0.35$ $4.06\pm 0.34$ $\mathbf{4.07}\pm\mathbf{0.43}$ $\mathbf{4.65}\pm\mathbf{0.32}$
Op-Sum-Gen $3.34\pm 0.68$ $3.92\pm 0.79$ $3.70\pm 0.54$ $\mathbf{4.0}\pm\mathbf{0.50}$ $\mathbf{4.08}\pm\mathbf{0.72}$ $3.87\pm 1.08$ $4.05\pm 1.31$
Ground-Truth $3.55\pm 0.50$ $3.93\pm 0.46$ $3.56\pm 0.31$ $4.08\pm 0.32$ $4.04\pm 0.46$ $3.81\pm 0.86$ $4.40\pm 0.45$
Table 19: Intrisic Evaluation results on the Amazon benchmark for all the models. Legend: AC: aspect-coverage, OPF: opinion-faithfulness, OPC: opinion-coverage, CC: conciseness, RE: relevance, HL: hallucination, LC: language-correctness.
$\cdot$ AC OPF OPC CC RL HL LC
Supervised $3.61\pm 0.22$ $4.10\pm 0.39$ $3.84\pm 0.33$ $4.04\pm 0.28$ $4.21\pm 0.31$ $4.19\pm 0.42$ $4.53\pm 0.27$
NaiveMean $3.56\pm 0.21$ $4.13\pm 0.41$ $3.84\pm 0.34$ $4.0\pm 0.32$ $\mathbf{4.31}\pm\mathbf{0.36}$ $4.26\pm 0.34$ $4.54\pm 0.39$
Synth-Feedback $3.56\pm 0.25$ $4.09\pm 0.40$ $3.79\pm 0.32$ $\mathbf{4.02}\pm\mathbf{0.30}$ $4.19\pm 0.34$ $4.19\pm 0.36$ $4.53\pm 0.29$
Inductive-Bias $\mathbf{3.63}\pm\mathbf{0.20}$ $\mathbf{4.22}\pm\mathbf{0.39}$ $\mathbf{3.85}\pm\mathbf{0.30}$ $4.01\pm 0.28$ $4.26\pm 0.29$ $\mathbf{4.33}\pm\mathbf{0.45}$ $\mathbf{4.61}\pm\mathbf{0.29}$
Ground-Truth $3.59\pm 0.15$ $3.88\pm 0.53$ $3.68\pm 0.27$ $4.02\pm 0.28$ $3.87\pm 0.59$ $3.67\pm 0.78$ $4.35\pm 0.44$
Table 20: Intrisic Evaluation results on the Flipkart benchmark for all the models. Legend: AC: aspect-coverage, OPF: opinion-faithfulness, OPC: opinion-coverage, CC: conciseness, RE: relevance, HL: hallucination, LC: language-correctness.
$\cdot$ AC OPF OPC CC RL HL LC
Supervised $3.47\pm 0.14$ $3.38\pm 0.26$ $3.49\pm 0.06$ $\mathbf{3.64}\pm\mathbf{0.19}$ $\mathbf{3.81}\pm\mathbf{0.26}$ $3.22\pm 0.56$ $3.96\pm 0.32$
NaiveMean $3.49\pm 0.05$ $3.48\pm 0.06$ $3.5\pm 0.0$ $3.56\pm 0.13$ $3.66\pm 0.22$ $3.52\pm 0.33$ $4.1\pm 0.33$
Synth-Feedback $3.50\pm 0.03$ $3.41\pm 0.26$ $3.5\pm 0.0$ $3.63\pm 0.24$ $3.62\pm 0.20$ $3.32\pm 0.63$ $\mathbf{4.03}\pm\mathbf{0.38}$
Inductive-Bias $\mathbf{3.54}\pm\mathbf{0.22}$ $\mathbf{3.50}\pm\mathbf{0.06}$ $\mathbf{3.57}\pm\mathbf{0.06}$ $3.62\pm 0.19$ $3.65\pm 0.23$ $\mathbf{3.68}\pm\mathbf{0.36}$ $4.0\pm 0.29$
Op-Sum-Gen $3.39\pm 0.3$ $3.46\pm 0.45$ $3.49\pm 0.28$ $3.61\pm 0.40$ $3.58\pm 0.82$ $3.43\pm 0.92$ $3.79\pm 1.18$
Ground-Truth $3.42\pm 0.22$ $3.475\pm 0.28$ $3.5\pm 0.0$ $3.57\pm 0.16$ $3.49\pm 0.28$ $3.21\pm 0.48$ $3.56\pm 0.23$
Table 21: Intrisic Evaluation results on the Oposum+ benchmark for all the models. Legend: AC: aspect-coverage, OPF: opinion-faithfulness, OPC: opinion-coverage, CC: conciseness, RE: relevance, HL: hallucination, LC: language-correctness.